### v1

# Existence and Uniqueness of Solutions for Steady Triple-Deck Equations

Abstract.
Triple-deck structures arise in boundary-layer flows when small localized wall perturbations induce moderate mean-flow distortion. Most existing studies in this field focus on the derivations of the asymptotic and numerical solutions of this system, while rigorous mathematical results remain scarce. In this paper, we establish, for the first time, the existence and uniqueness of the solutions to the steady triple-deck equations. The steady triple-deck equations can be reduced to an Airy equation with frequency as a parameter via Fourier transformation. The main difficulty is that the Airy equation degenerates at zero frequency, hindering uniform estimates of the solution with respect to the frequency parameter. A key contribution is that we find a subtle function involving the distribution of zeros of the Bessel function to overcome the degeneracy by analyzing the structure of the Airy function. Additionally, we derive a higher-frequency estimate and modify an elliptic multiplier to obtain uniform estimates of the solution.
1. Introduction
In 1904, Prandtl [21] proposed the boundary layer theory to address attached flow problems, achieving remarkable success in fluid dynamics research. However, when dealing with flow separation induced by adverse pressure gradients, this theory encounters singularity issues. Notably, Landau and Lifshitz [14] revealed that the velocity component normal to the body surface grows unboundedly (termed the Goldstein singularity), which hinders the solution from being further extended to the reverse-flow region [6]. To overcome these difficulties, scholars including A.F. Messiter [17] and K. Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s. In this theory, three distinct layers—a viscous sublayer, a main layer, and an upper layer—appear in the transverse direction, for which uniformly valid approximate solutions are derived by the asymptotic matched method. For instance, Messiter successfully eliminated the Goldstein singularity in his study on the flow near the trailing edge of an incompressible flat plate, demonstrating the theory’s mathematical resolution to singular behaviors [17].
In physics, the triple-deck structure is observed when localized surface perturbations exist. For demonstration, we focus on the steady triple-deck flow over an incompressible two-dimensional (2D) flat plate with a localized roughness element. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and we denote it as $\{(X,Y)\big{|}\ Y=\epsilon^{5}F(\frac{X-1}{\epsilon^{3}}),\ X\in{\mathbb{R}}\}$ for some given function $F=O(1)$ as $\epsilon\to 0$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (see figure 1).
Figure 1. Demonstration of the triple-deck structure (not to scale): I Viscous Sublayer; II Main Layer; III Upper Layer; $\delta$ Thickness of Prandtl layer.
We denote the velocity vector by ${\mathbf{u}}_{NS}=(u_{NS}(X,Y),v_{NS}(X,Y))$, the pressure $p_{NS}$, and demonstrate their scales in Table 1.
I II III
$u_{NS}$ $O(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+O(\epsilon)$ $1+O(\epsilon^{2})$
$v_{NS}$ $O(\epsilon^{3})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
$p_{NS}$ $O(\epsilon^{2})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
Table 1. Scales of velocities and pressure in the triple-deck structure: $\tilde{U}_{0}$ Streamwise velocity in Prandtl layer.
It is well-known that the 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
(1.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
Based on the scaling relations in Figure 1 and Table 1, and by applying the method of matched asymptotic expansions (see details in Appendix), we derive that the flows in the upper layer and main layer are given by
(1.2) $u_{NS}\sim 1-\epsilon\Re(f(z)),\quad v_{NS}\sim\epsilon^{2}\Im(f(z)),\quad p_{NS}\sim\epsilon^{2}\Re(f(z)),\quad z=x_{*}+i\bar{y},$
$u_{NS}\sim\tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad v_{NS}\sim-\epsilon^{2}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad p_{NS}\sim\epsilon^{2}\tilde{p}(x_{*}).$
Here $x_{*}=\frac{X-1}{\epsilon^{3}},\ \bar{y}=\frac{Y}{\epsilon^{3}},\ \tilde{y}=\frac{Y}{\epsilon^{4}}$, $\tilde{U}_{0}^{\prime}(0)=1$, $A_{*}$ is an unknown function (called the displacement of the velocity), and
(1.3) $f(z)=-\frac{1}{\pi}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}}{\zeta-z}d\zeta,\quad\tilde{p}(x_{*})=\lim_{\Im(z)\to 0^{+}}\Re(f(z))=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
For the viscous sublayer (region I), we derive the governing equations as
(1.4) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\end{aligned}\right$.
Here $y_{*}=\frac{Y}{\epsilon^{5}}$ and $(\epsilon u_{*},\epsilon^{3}v_{*},\epsilon^{2}P_{*})$ are the leading-order terms of $(u_{NS},v_{NS},p_{NS})$. The velocity field is subject to the no-slip, non-penetration conditions,
(1.5) $(u_{*},v_{*})\big{|}_{y_{*}=F(x_{*})}=0.$
In the upstream limit, the velocity approaches the unperturbed Blasius profile,
(1.6) $(u_{*}-y_{*})\big{|}_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big{|}_{x_{*}\downarrow-\infty}=0.$
To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Since the Blasius solution satisfies $\tilde{U}_{0}\to\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$, then for any $\beta_{0}>0$, it follows that
(1.7) $u_{NS}=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y})+o(\epsilon)=\tilde{y}+\epsilon A_{*}(x_{*})+o(\epsilon+\tilde{y}^{4-\beta_{0}})$
$=$ $\ \epsilon\big{(}y_{*}+A_{*}(x_{*})\big{)}+o\big{(}\epsilon+\epsilon^{(4-\beta_{0})(1-\alpha_{0})}\big{)}.$
To ensure the second part to be negligible, we require $(4-\beta_{0})(1-\alpha_{0})\geq 1$, namely, $\alpha_{0}\in(0,\frac{3}{4})$ and $\beta_{0}\in(0,4)$. Likewise, we can derive that the leading-order pressure in the sublayer and that in the main layer are equal, that is,
(1.8) $P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
In light of (1.7), it is permissible to set the upper boundary of the viscous sublyaer domain to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$, where $f$ is to be determined. Consequently, the upper boundary condition is
(1.9) $(u_{*}(x_{*},y_{*})-y_{*})\big{|}_{y_{*}=f(x_{*})}=A_{*}(x_{*}).$
By (1.4)-(1.6), and (1.9), we derive the closed system for the viscous sublayer as
(1.10) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &(u_{*},v_{*})\big{|}_{y_{*}=F(x_{*})}=0,\quad(u_{*}(x_{*},y_{*})-y_{*})\big{|}_{y_{*}=f(x_{*})}=A_{*}(x_{*}),\quad x_{*}\in{\mathbb{R}},\\ &(u_{*}-y_{*})\big{|}_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big{|}_{x_{*}\downarrow-\infty}=0,\quad F(-\infty)<y_{*}<f(-\infty),\\ &P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.\end{aligned}\right$.
A critical observation is that the flows (1.2) in the upper two layers only depend on the unknown quantities $(u_{*},v_{*},A_{*},P_{*})$ within the viscous sublayer, this paper thus focuses on the equations (1.10) of the viscous sublayer (henceforth referred to as the triple-deck equations).
1.1. Known results
Mathematically, the triple-deck equations share a structural resemblance to the Prandtl equations of classical boundary layers, yet they differ fundamentally in domain and pressure determination. The Prandtl equations are defined over $(0,L_{0})\times\mathbb{R}_{+}$, with pressure $P_{*}$ dictated by external potential flow $u_{E}(x_{*})$ via the Bernoulli relation
(1.11) $P_{*}+\frac{1}{2}u_{E}^{2}=\text{Const,}$
which means that the pressure is a given function. There are a lot of results on this topic. The global existence of solutions to the Prandtl equations was first established by Oleinik [19], and subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30]. Moreover, the flow separation and reverse phenomena have also attracted the attention of many scholars [2, 10, 15, 25].
Compared with (1.11), the pressure of triple-deck model is given by (1.8) which is defined by unknown function $A_{*}$. Due to that, to construct a solution of triple-deck model is not an easy way. For the unsteady case, Iyer and Vicol [11], Gérard-Varet et al. [5] proved the well-posedness of solutions in real analytic spaces, while Dietert and Gérard-Varet [3] revealed instability for some smooth initial data, indicating that the class of analytic functions is optimal for well-posedness. For steady cases, mathematical results are sparser. When $F=0$, Iyer and Maekawa [9] proved a local Liouville-type theorem, pointing out that under specific rigid conditions, the equations admit only the Couette solution. However, physicists found that the triple-deck structure only appears in flow with some perturbations which means that $F\not\equiv 0$. For more results on physical analysis and numerical simulations, see [1, 4, 12, 26, 32]. It is noteworthy that, for the triple-deck equations (1.10) in the presence of disturbance, the well-posedness of solutions lacks rigorous mathematical conclusions. This paper investigates the triple-deck equations (1.10) for some $f=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,\frac{3}{4})$, establishes the existence and uniqueness theory in the presence of disturbance for the first time, and verifies the compatibility between the disturbance scale and physical observation results.
1.2. Main results
To simplify the equations (1.10), we introduce the new coordinate $(x,y)$ and new notations $(u,v,P,A)$:
(1.12) $x=x_{*},\ y=y_{*}-F(x_{*}),$
$(u_{*},v_{*},P_{*},A_{*})=(y+u,v+(y+u)\frac{dF}{dx},P,A).$
Then, the problem (1.10) can be reduced to
(1.13) $\left\{\begin{aligned} &y\partial_{x}u+v+\partial_{x}|\partial_{x}|A-\partial_{y}^{2}u=-u\partial_{x}u-v\partial_{y}u,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{x}u+\partial_{y}v=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &(u,v)\big{|}_{y=0}=(0,0),\quad u\big{|}_{y=f(x)-F(x)}=A(x)+F(x),\quad x\in{\mathbb{R}},\\ &\lim_{x\to-\infty}u=0,\quad 0<y<f(-\infty)-F(-\infty).\end{aligned}\right$.
For convenience, we take $f(x)=M+F(x)$ for $M=\epsilon^{-\alpha_{0}}$ and $\alpha_{0}\in(0,\frac{3}{4})$. It is useful to introduce the vorticity $\omega=\partial_{y}u$ and
(1.14) $u(x,y)=I_{y}[\omega]:=\int_{0}^{y}\omega dy^{\prime}.$
Differentiating the first equation of (1.13) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows.
(1.15) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Here $F$ is a prescribed wall disturbance function. We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$.
To establish a rigorous framework, we introduce some functional spaces in this paper. Define the weighted Sobolev spaces for vorticity $\omega$ as
(1.16) $X_{\bar{\alpha}}=$ $\ \bigg{\{}f\big{|}\ {\big{\|}f\big{\|}}_{X_{\bar{\alpha}}}={\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}<\infty\bigg{\}},$
$Y_{\bar{\alpha}}=$ $\ \bigg{\{}f\big{|}\ {\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}={\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\qquad\qquad\quad+{\big{\|}|\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}f\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\qquad\qquad\quad+{\big{\|}|\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}<\infty\bigg{\}},$
where $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$.
For the displacement function $A$, we define
(1.17) $X_{\alpha,\infty}=\bigg{\{}f\big{|}\ {\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
The combined space for $(\omega,A)$ incorporates the upper boundary condition is
(1.18) ${\mathcal{X}}_{\alpha,M}=\bigg{\{}(\omega,A)\big{|}$ $\ {\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}={\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A\big{\|}}_{X_{\alpha,\infty}}<\infty,$
$\ \ \omega(\cdot,y)\big{|}_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg{\}}.$
Here $Ai(z)$ is the Airy function satisfying $\frac{d^{2}}{dz^{2}}Ai(z)-zAi(z)=0,\ Ai\big{|}_{z\uparrow+\infty}=0$.
We adopt standard mixed-norm notations on $\Omega_{M}$ as following:
(1.19) $\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big{\|}f\big{\|}}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
${\big{\|}f\big{\|}}_{L_{x}^{p}L_{y}^{q}}={\big{\|}f\big{\|}}_{L_{x}^{p}({\mathbb{R}};L_{y}^{q}([0,M]))},\quad\mbox{for }\ p,q\geq 1,$
${\big{\|}f\big{\|}}_{H_{x}^{s}L_{y}^{2}}={\big{\|}f\big{\|}}_{H_{x}^{s}({\mathbb{R}};L_{y}^{2}([0,M]))},\quad{\big{\|}f\big{\|}}_{L_{x}^{2}H_{y}^{s}}={\big{\|}f\big{\|}}_{L_{x}^{2}({\mathbb{R}};H_{y}^{s}([0,M]))},\quad\mbox{for }\ s\in{\mathbb{R}}.$
For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$, such that $B_{0}\leq CB_{1}$.
To seek the solution of the system (1.15), we use the following decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ where $(\omega_{0},A_{0})$ satisfies the following linear part of (1.15):
(1.20) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Our first result is to construct a solution of (1.20).
Theorem 1.1.
Let $\alpha\in(0,\frac{7}{3}]$, then the linearized system (1.20) has a solutions $(\omega_{0},A_{0})$ belongs to ${\mathcal{X}}_{\alpha,M}$ and satisfies the following estimates:
(1.21) ${\big{\|}\omega_{0}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A_{0}\big{\|}}_{X_{\alpha,\infty}}\leq$ $\ C{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})},$
where $C>0$ is independent of $\alpha,M$, and $F$.
Now, we are the position to state the main results of this paper.
Theorem 1.2.
Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$. Then, there exist a small enough constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and large enough constant $M_{0}$, such that when $M\geq M_{0}$ and
${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\delta_{0}M^{-\frac{5+3\varepsilon}{2}},$
the system (1.15) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ where $(\omega_{0},A_{0})$ defined by (1.20) and $(\bar{\omega},\bar{A})$ satisfying
(1.22) ${\big{\|}\bar{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}\big{\|}}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2},$
where $C>0$ is independent of $\alpha,\varepsilon,M$, and $F$. Moreover, $(\omega,A)$ is the unique solution of (1.15) within ${\mathcal{X}}_{\alpha,M}$.
Remark 1.1.
Our theoretical results align closely with physical phenomena. The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$ under small perturbations $F$, consistent with the principle that linear approximations govern near-equilibrium states in fluid dynamics.
Remark 1.2.
Considering the scaling relation $M=O(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$, our analysis shows that the allowable perturbation scale $O(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
Remark 1.3.
When $F=0$, our results imply that there is a unique zero solution in ${\mathcal{X}}_{\alpha,M}$ which aligns with the results by Iyer and Maekawa in [9].
Remark 1.4.
Here, we take $M\sim\epsilon^{-\alpha_{0}}$ with some small data $\epsilon$ which implies that $M$ will be large enough. Thus, taking $M_{0}$ large enough is satisfied by our model. On the other hand, for general $M_{0}$, the results may also be got by Remark 3.2.
1.3. Main ideas
In this subsection, we state the main ideas of the paper. First, we notice that the system (1.15) is a degenerate nonlinear parabolic equation with variable coefficients, where the degeneration occurs on the wall $y=0$. To deal with the degeneration, we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ satisfying
(1.23) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}+F,\quad x\in{\mathbb{R}},\end{aligned}\right$.
and
(1.24) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(u\partial_{x}\omega+v\partial_{y}\omega)\triangleq{\bar{h}},\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\bar{\omega}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A},\quad x\in{\mathbb{R}}.\end{aligned}\right$.
It separates the main effects caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
In the homogeneous problem (1.23), we observe that the integral condition exhibits nonlocal behavior, so it cannot be treated in isolation. Instead, a unified framework is needed to couple this integral condition with the boundary data at $y=0$ (i.e., $\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0}$). Their synthesis yields an elliptic multiplier $m(D)^{-1}$ (see (3.45)). This multiplier crucially links the displacement $A_{0}$ and vorticity $\omega_{0}$ by
(1.25) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\lesssim{\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{L^{2}}\lesssim{\big{\|}F\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})},\ \bar{\alpha}\geq 0,$
providing a key tool for analyzing the well-posedness of the pair $(\omega,A)$.
In order to solve the non-homogeneous problem (1.24), the nonlinear terms and boundary terms are treated separately. Then, we decompose $\bar{\omega}=\bar{\omega}_{e}+\bar{\omega}_{b}$ satisfying
(1.26) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e}-\partial_{y}^{2}\bar{\omega}_{e}=-(u\partial_{x}\omega+v\partial_{y}\omega),\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{e}\big{|}_{y=0}=0,\quad x\in{\mathbb{R}},\end{aligned}\right$.
and
(1.27) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b}-\partial_{y}^{2}\bar{\omega}_{b}=0,\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{b}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}_{b}]=\bar{A}-I_{M}[\bar{\omega}_{e}],\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Similar to the homogeneous problem (1.23), we can find a solution $(\bar{\omega}_{b},\bar{A})$ to (1.27) with the following estimates
(1.28) ${\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{\bar{\alpha}}}\lesssim$ $\ {\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}\bar{A}\big{\|}}_{L^{2}}\lesssim{\big{\|}I_{M}[\bar{\omega}_{e}]\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})}\quad\text{for }\ \bar{\alpha}\geq 0$
[…]
Hence, the essential step is to solve (1.26). To this end, we firstly omit the integral condition and consider a linear problem as following
(1.29) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big{|}_{y=0}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Upon Fourier transformation, the linearized operator of the first equation in (1.29) is expressed as $i\xi y-\partial_{y}^{2}$, and it shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$. Then it is straight to observe that there exists a solution $\hat{f}_{e}$ to the following linearized problem
$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big{|}_{y=0}=0,\quad\xi\in{\mathbb{R}},\end{aligned}\right$.
for any given function $h\in L^{2}(\Omega_{M})$, and $\hat{f}$ is the Fourier transformation of $f$ in the variable $x$. Moreover, similar to the proof in [9], the solution satisfies
(1.30) ${\big{\|}f_{e}\big{\|}}_{Y_{0}}\leq C{\big{\|}h\big{\|}}_{L^{2}}.$
Here $C>0$ is independent of $M$. Thus, in order to solve (1.26), we need to obtain the $L^{2}$-estimate of the nonlinear terms $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$. By calculations, we obtain
(1.31) ${\big{\|}u\partial_{x}\omega+v\partial_{y}\omega\big{\|}}_{L^{2}}\lesssim({\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}+{\big{\|}y^{\frac{1}{2}}\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}+{\big{\|}y\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}^{2}$
$\qquad\qquad\qquad\qquad\quad+{\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data}.$
Clearly, the norms ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}$ and ${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$ can not be controlled by ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}$. Even if we add low-frequency information, the subspace $L^{2}(\Omega_{M})\cap L_{x}^{2}\dot{H}_{y}^{2}\cap\dot{H}_{x}^{\frac{2}{3}}L_{y}^{2}$ of $Y_{0}$ is unable to be embedded into $L^{\infty}$. Thus, we need not only a low-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}(\Omega)}$, but also a higher-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}$ for $\alpha>0$ such that
(1.32) ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{\alpha}}^{1-\theta_{1}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{\alpha}}^{1-\theta_{1}},\ \text{for some }\theta_{1}\in(0,1),$
${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}^{\theta_{2}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}^{1-\theta_{2}},\ \text{for some }\theta_{2}\in(0,1).$
The main difficulty of this paper is to obtain the low-frequency estimate of $\bar{\omega}_{e}$. We observe that the homogeneous solution of the linearized operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$. For any nonhomogeneous term $h$, the general solution includes a term of the form $(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz$, where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function. Note that the nonlinear term $u\partial_{x}\omega-v\partial_{y}\omega$ cannot be entirely recast solely as $x$-directional derivatives (i.e., $\partial_{x}(\text{Terms})$). This structural constraint hinders low-frequency estimates derived from the nonlinear terms. To address the singularity near $\xi=0$, a key observation is that the function
$t(z)=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}$
has no zeros in the sector $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})\}$, where $Ai(z),Bi(z)$ denote two independent Airy functions. This property is non-trivial (see Remark 3.1), which involves the distribution of zeros of the Bessel function. Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form:
(1.33) $G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
$\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y),$
$\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
It is straight to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\}$. With these observations, we derive that
(1.34) ${\big{\|}f_{e}\big{\|}}_{L^{2}}\lesssim{\big{\|}h\big{\|}}_{L^{2}}.$
To control the higher-frequency of $f_{e}$, we introduce the space $Y_{\alpha}$ for $\alpha>0$. It requires higher regularity of $h$ since
(1.35) ${\big{\|}f_{e}\big{\|}}_{Y_{\alpha}}\lesssim{\big{\|}h\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}},\quad\forall\alpha>0.$
By product estimates, we finally control the $H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$-norm of nonlinear terms in $\bar{h}$} for $\alpha>\frac{2}{3}$, that is,
${\big{\|}{\bar{h}}\big{\|}}_{H_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},\quad\forall\alpha>\frac{2}{3}.$
It follows that
(1.36) ${\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data},\quad\forall\alpha>\frac{2}{3}.$
Although the above process is sufficient to obtain a solution $(\omega,A)$ in $L^{2}\cap Y_{0}\cap Y_{\alpha}$ for $\alpha>\frac{2}{3}$, it is natural to require that the displacement $A$ is bounded for all $M\geq M_{0}$. We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then it satisfies
${\big{\|}A\big{\|}}_{L^{\infty}({\mathbb{R}})}\leq{\big{\|}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}.$
Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ estimate of $\omega$ via the norm ${\big{\|}\omega\big{\|}}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$. To this end, we utilize the interpolation inequality. Let $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$ be an interpolation exponent, then we have
${\big{\|}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq C(\bar{\alpha},M){\big{\|}\omega\big{\|}}_{L^{2}}^{\theta_{1}}{\big{\|}\omega\big{\|}}_{Y_{0}}^{1-\theta_{1}}+{\big{\|}\omega\big{\|}}_{L_{x}^{\infty}({\mathbb{R}};L_{y}^{1}([0,1])}\leq C_{0}(\bar{\alpha},M){\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}.$
Here $\bar{\alpha}\in[0,\frac{7}{3}]$ and $M\geq M_{0}$. We observe that
(1.37) $\inf_{\bar{\alpha}\in[0,\frac{7}{3}]}C_{0}(\bar{\alpha},M)=C_{0}(0,M)=O([\log(M)]^{\frac{1}{2}}),\quad\forall M\geq M_{0}.$
It implies that the bound of $A$ can not be controlled uniformly in $M$ by the norm ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}$, which contradicts the physical observations that the linear part $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ determines the behavior of $(\omega,A)$. To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big{\|}}_{L^{2}}$, so as to derive a refined estimate. Specifically, we obtain
${\big{\|}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big{\|}}_{L^{2}}+{\big{\|}\omega\big{\|}}_{Y_{0}}\big{)}\leq C{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}.$
Here $C>0$ is independent of $M$. And the estimate of the new norm is derived by using more information of Airy functions (see Proposition 2.3).
1.4. Organization of the paper
This paper is organized as follows. First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key a priori estimates, including weighted Sobolev norms for vorticity and displacement. Subsequently, Section 3 focuses on the linearized problem of the triple-deck system. By decomposing the linearized equations into well-posed subproblems (e.g., boundary-value problems for the Airy equation), we construct solutions to the linearized system and establish quantitative bounds on their regularity. Thereafter, in Section 4, we apply these linear estimates to prove the main result (Theorem 1.2) via a fixed-point iteration scheme, establishing the existence and uniqueness of solutions to the full nonlinear triple-deck equations. Finally, we present an appendix, which provides a rigorous formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer. We also present explicit solutions for the upper and main layers, demonstrating their dependence on the viscous sublayer solution. This appendix reinforces the self-contained nature of the mathematical framework.
2. Preliminaries
To study the quantity $\omega$, we first consider the space $X_{0}\cap X_{\alpha}$. Note that the space $X_{\bar{\alpha}}$ consists of a low-frequency part and a high-frequency part. The low-frequency part is characterized by the norms ${\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}$ and ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}$, while the high-frequency part is characterized by ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}$. We can explore various properties between these two parts. First, we present some of them for $f\in X_{0}$.
Proposition 2.1.
Assume that $f\in X_{0}$, for any $0\leq\varepsilon\leq\frac{1}{6}$, we have
(2.1) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}+{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}+{\big{\|}f\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big{\|}f\big{\|}}_{X_{0}}.$
Here $C$ is independent of $\varepsilon$ and $f$.
Proof.
By Sobelev embedding theorem and the interpolation theorem, it is straight to obtain
(2.2) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}H_{y}^{2}}\leq C{\big{\|}f\big{\|}}_{X_{0}},$
${\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}\leq{\big{\|}|\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{6\varepsilon}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{1-6\varepsilon}\leq{\big{\|}f\big{\|}}_{Y_{0}}.$
Using the inequality ${\big{\|}|\partial_{x}|^{-\frac{1}{3}}f\big{\|}}_{L_{x}^{6}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}}$, we derive that
(2.3) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}.$
Having established the above estimates, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through some calculations, we obtain
(2.4) ${\big{\|}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg{\{}{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big{\|}y\partial_{x}f\big{\|}}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
$\qquad\qquad\qquad\qquad\quad\ {\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{3}{11}}{\big{\|}|\partial_{x}|^{\frac{2}{3}}f\big{\|}}_{L_{x}^{2}}^{\frac{8}{11}}\cdot y^{-\frac{1}{22}}\bigg{\}},$
${\big{\|}f\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{3}{11}}{\big{\|}|\partial_{x}|^{\frac{2}{3}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{8}{11}}{\big{\|}y^{-\frac{1}{22}}\big{\|}}_{L_{y}^{2}([0,1])}$
$\qquad\qquad\qquad+C{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{9}{17}}{\big{\|}y\partial_{x}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{8}{17}}{\big{\|}y^{-\frac{19}{34}}\big{\|}}_{L_{y}^{2}([1,\infty])}$
$\qquad\qquad\leq C{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{3}{11}}{\big{\|}|\partial_{x}|^{\frac{2}{3}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{8}{11}}$
$\qquad\qquad\qquad+C{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{9}{17}}{\big{\|}y\partial_{x}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{8}{17}}$
$\qquad\qquad\leq C{\big{\|}f\big{\|}}_{X_{0}}.$
Observe that as the index $\alpha$ increases, higher frequencies in the variable $x$ is accounted for in $X_{\alpha}$, since ${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}f\big{\|}}_{Y_{0}}={\big{\|}f\big{\|}}_{Y_{\alpha}}$. These higher frequencies lead to other useful properties.
Proposition 2.2.
Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$, then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we obtain
(2.5) ${\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}+{\big{\|}f\big{\|}}_{L^{\infty}(\Omega_{M})}\leq C\bigg{(}\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\bigg{)}({\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}).$
Here $C>0$ is independent of $\varepsilon,\alpha$, and $f$.
Proof.
By interpolation and Proposition 2.1, we have
(2.6) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big{\|}|\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
${\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}\leq\frac{C}{\sqrt{\varepsilon}}{\big{\|}|\partial_{x}|^{\frac{1}{2}+\varepsilon}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\bigg{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\bigg{)}{\big{\|}f\big{\|}}_{Y_{0}}^{\frac{1}{2}}$
$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}({\big{\|}f\big{\|}}_{Y_{0}}+{\big{\|}f\big{\|}}_{Y_{\alpha}}).$
By classical extension, we denote $\tilde{f}$ as the extension of $f$ to ${\mathbb{R}}^{2}$. It is evident to verify that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
(2.7) ${\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
${\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\leq C{\big{\|}f\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}(0,M))}.$
By calculations, we have
(2.8) ${\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\hat{\tilde{f}}\big{\|}}_{L_{\xi,\eta}^{1}}$
$=$ $\ \int_{{\mathbb{R}}^{2}}\bigg{\{}\big{|}\hat{\tilde{f}}(\xi,\eta)(1+|\eta|)^{2}\big{|}^{\theta}\big{|}\hat{\tilde{f}}(\xi,\eta)(1+|\xi|)^{\frac{2}{3}+\frac{\alpha}{2}}\big{|}^{1-\theta}$
$\qquad\quad\cdot(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\bigg{\}}d\xi d\eta$
$\leq$ $\ C(\alpha,\theta){\big{\|}\big{|}\hat{\tilde{f}}(\xi,\eta)(1+|\eta|)^{2}\big{|}^{\theta}\big{\|}}_{L_{\xi,\eta}^{\frac{2}{\theta}}}{\big{\|}\big{|}\hat{\tilde{f}}(\xi,\eta)(1+|\xi|)^{\frac{2}{3}+\frac{\alpha}{2}}\big{|}^{1-\theta}\big{\|}}_{L_{\xi,\eta}^{\frac{2}{1-\theta}}}$
$\leq$ $\ C(\alpha,\theta){\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}^{\theta}{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}^{1-\theta}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}.$
Here $C(\alpha,\theta)={\big{\|}(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big{\|}}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$ in the last inequality. It implies that
(2.9) ${\big{\|}f\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big{\|}f\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big{(}{\mathbb{R}};L_{y}^{2}((0,M))\big{)}}\bigg{)}$
$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}\bigg{)}.$
Next, we present some a priori estimates for the nonlinear terms. Suppose $(u,v)$ satisfies
$\partial_{x}u+\partial_{y}v=0,$
with the boundary data
$(u,v)\big{|}_{y=0}=0,\quad u\big{|}_{y=M}=A_{M}.$
Set $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$ for $\omega=\partial_{y}u$, then we have
(2.10) $h=-u\partial_{x}\omega-\bigg{(}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz-y\partial_{x}A_{M}\bigg{)}\partial_{y}\omega.$
Define
(2.11) $V^{0}=\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz.$
Next, we give the estimates for the nonlinear term $h$.
Proposition 2.3.
Fixing $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$, suppose $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$, then we have $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$. Moreover, there holds
(2.12) ${\big{\|}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},\quad\forall\ \varepsilon\in(0,\frac{1}{6}]$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Here $C=C(\alpha,\varepsilon)>0$ is independent of $M,\ \omega$, and $A_{M}$.
Proof.
In view of Proposition 2.1 and Proposition 2.2, it is straight to obtain
${\big{\|}h\big{\|}}_{L^{2}(\Omega_{M})}\leq$ $\ C{\big{\|}\frac{u}{y}\big{\|}}_{L^{\infty}(\Omega_{M})}{\big{\|}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}+CM^{\frac{3\varepsilon}{2}}{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
$\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
$\leq$ $\ C({\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}$
$\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
which implies the first desired result.
For the high-order estimates, by product estimates, we have
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})},$
and
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}$
$\qquad\qquad+C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{4}L_{y}^{\infty}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
$\qquad\qquad+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}.$
All we left is to give the estimates of the right hand of above inequality. By Sobolev embedding Theorem, we have
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+C{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{3}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\leq C{\big{\|}\omega\big{\|}}_{X_{\alpha}},$
${\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}\leq\ M{\big{\|}|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega)}$
$\qquad\leq C\frac{[1+(\alpha-\frac{2}{3})]^{\frac{1}{2}}}{\alpha-\frac{2}{3}}M{\big{\|}|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}+(\frac{\alpha}{2}-\frac{1}{3})}L_{y}^{2}\cap L_{x}^{2}H_{y}^{2}}\quad\mbox{for}\ \ \alpha>\frac{2}{3}$
$\qquad\leq CM\big{(}{\big{\|}|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{1}{3}}\partial_{y}^{2}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}$
$\qquad\leq CM\big{(}{\big{\|}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}\partial_{y}^{2}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{y}^{2}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}$
$\qquad\leq CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},$
and
${\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq C\big{(}{\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}\leq{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{4}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}$
$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Combining all above estimates, for $\alpha>\frac{2}{3}$, we obtain
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}\leq$ $\ CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big{\|}}_{L^{2}(\Omega_{M})}\leq$ $\ CM^{\frac{3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}.$
Moreover, we have
(2.13) ${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(y\partial_{x}A_{M}\partial_{y}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}y\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}+C{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
$\qquad\leq CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}}$
$\qquad\leq CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
Remark 2.1.
Indeed, if $h$ is replaced by $h_{1}=-(u\partial_{x}\omega_{1}+v\partial_{y}\omega_{1})$ for $\omega_{1}\in X_{0}\cap X_{\alpha}$, then by repeating the proof of Proposition 2.3, we have
(2.14) ${\big{\|}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}},\ \ \forall\varepsilon\in(0,\frac{1}{6}]$
${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Next we give priori estimate of $I_{M}[f]$ for $f\in X_{\bar{\alpha}}$.
Proposition 2.4.
Suppose $\bar{\alpha}\geq 0$, for any $f\in X_{\bar{\alpha}}$, we have
(2.15) ${\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{\|}}_{L^{2}({\mathbb{R}})}\leq$ $\ C{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}.$
Here $C$ is independent of $\bar{\alpha}$ and $f$.
Proof.
Observe that
$\big{|}\int_{0}^{M}\hat{f}(\xi,y)dy\big{|}=$ $\ \big{|}\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\big{|}$
$\leq$ $\ C|\xi|^{-\frac{1}{6}}{\big{\|}\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}}+C|\xi|^{\frac{1}{6}}{\big{\|}y\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}},$
which implies that
(2.16) $\big{|}\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{|}\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big{\|}\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big{\|}y\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}}).$
Thus,
(2.17) ${\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{\|}}_{L^{2}({\mathbb{R}})}\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L^{2}}+{\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big{\|}}_{L^{2}}\big{)}\leq C{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}},$
which implies the desired results. ∎
3. Proof of Theorem 1.1
In this section, we give the proof of the Theorem 1.1. Before that, we study a nonhomogeneous linearized problem without the integral condition as follows.
(3.1) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h_{e},\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big{|}_{y=0}=0,\quad f_{e}\big{|}_{y=M}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Lemma 3.1.
Fix $\alpha>0$. We assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$. Then, there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.1). Moreover, there holds
${\big{\|}f_{e}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C\big{[}(1+M^{2}){\big{\|}h_{e}\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big{\|}}_{L^{2}(\Omega_{M})}\big{]},$
where $C$ is a constant independent of $\alpha,M$, and $h_{e}$.
Proof.
The proof proceeds in three steps.
Step 1. We give the explicit form of the solution. Taking the Fourier transform in variable $x$ into (3.1), we obtain
(3.2) $\left\{\begin{aligned} &i\xi y\hat{f}_{e}-\partial_{y}^{2}\hat{f}_{e}=\hat{h}_{e},\quad(\xi,y)\in{\mathbb{R}}\times(0,M),\\ &\partial_{y}\hat{f}_{e}\big{|}_{y=0}=0,\ \hat{f}_{e}\big{|}_{y=M}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.2), which has an explicit formula as following:
(3.3) $\hat{f}_{e}(\xi,y)=$ $\ (i\xi)^{-\frac{1}{3}}\int_{0}^{M}G(y,z;\xi)\hat{h}_{e}(\xi,z)dz,$
with
(3.4) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
$s_{M}=\bigg{(}\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg{)}^{-1},$
$\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y),$
$\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
Step 2. We derive the low-frequency estimates of $f_{e}$. Since $\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=-\frac{1}{\sqrt{3}}$, we have
(3.5) $t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
It is well-known that $Ai(z)$ and $Bi(z)$ are holomorphic on $\mathbb{C}$} and do not vanish for $\text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$ [28, Page 15]. For clarity of the proof, we claim that $\sqrt{3}Ai(z)+Bi(z)$ on $\mathbb{C}$} has all its zeros in the rays $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ \ k\in{\mathbb{Z}}\big{\}}$ (see details in Remark 3.1). Define a sector with the angle $2\theta$ by
$C_{\theta}=\big{\{}z\in{\mathbb{C}}\big{|}\ \text{Arg}(z)\in(-\theta,\theta)\big{\}}.$
Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$. Moreover, for any $\delta\in(0,\frac{\pi}{3})$, there holds
(3.6) $Ai(z)\sim\frac{1}{3^{\frac{2}{3}}\Gamma(\frac{1}{3})}-\frac{1}{3^{\frac{1}{3}}\Gamma(\frac{2}{3})}z,\quad Bi(z)\sim\frac{1}{3^{\frac{1}{6}}\Gamma(\frac{1}{3})}+\frac{3^{\frac{1}{6}}}{\Gamma(\frac{2}{3})}z,\qquad\mbox{as}\ \ z\to 0,$
$Ai(z)\sim\frac{e^{-\frac{2}{3}z^{\frac{3}{2}}}}{2\sqrt{\pi}z^{\frac{1}{4}}},\quad Bi(z)\sim\frac{e^{\frac{2}{3}z^{\frac{3}{2}}}}{\sqrt{\pi}z^{\frac{1}{4}}},\qquad\mbox{as}\ \ |z|\to\infty\ \ \big{(}z\in C_{\frac{\pi}{3}-\delta}\big{)}.$
By (3.6), we have
(3.7) $\lim_{|z|\to\infty,\atop|\text{Arg}(z)|\in[0,\frac{\pi}{3}-\delta)}t(z)=\frac{1}{\sqrt{3}}.$
Then, combining the holomorphicity and non-vanishing property of $t(z)$ on $C_{\frac{\pi}{3}}$ with (3.7), there exists some $c_{0}>0$ such that
$|t(z)|\geq c_{0}^{-1},\quad\text{for }z\in\overline{C_{\frac{\pi}{6}}}.$
Since $s_{M}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we have
(3.8) $|s_{M}|\leq c_{0},\qquad\forall\ \xi\in{\mathbb{R}},\ M>0.$
Then for $|\xi|\leq M^{-3}$, we obtain
(3.9) $|\mu_{M}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}$
$\qquad\qquad\ \leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M],$
$|\mu_{0}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}\leq C,\quad\forall\ y\in[0,M].$
It follows that
(3.10) $|\hat{f}_{e}(\xi,y)|\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
$\qquad\qquad\qquad\quad+\big{|}\mu_{0}(\xi,y)\big{|}\cdot\int_{y}^{M}\big{|}\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz\bigg{\}}$
$\leq$ $\ C(M-y)y^{\frac{1}{2}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([0,y])}$
$\quad+C|\xi|^{-\frac{1}{3}}{\big{\|}|\xi|^{\frac{1}{3}}(M-z)\big{\|}}_{L_{z}^{2}([y,M])}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([y,M])}$
$\leq$ $\ C(M-y)M^{\frac{1}{2}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
Hence, we have
(3.11) ${\big{\|}(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big{\|}}_{L_{y}^{2}(0,M)}\leq$ $\ CM^{2}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}(0,M)},$
for $\alpha^{\prime}\geq 0$.
And for $|\xi|\geq M^{-3}$, we have
(3.12) $|\mu_{M}(\xi,y)|=\big{|}Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\big{|}$
$\qquad\qquad\ \leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
$|\mu_{0}(\xi,y)|=\big{|}Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\big{|}$
$\qquad\qquad\leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}\bigg{[}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}+e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\bigg{]}$
$\qquad\qquad\leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
and
(3.13) ${\big{\|}\mu_{0}(\xi,z)\big{\|}}_{L_{z}^{2}(0,y)}\leq C|\xi|^{-\frac{1}{6}}\bigg{(}{\big{\|}Ai(z)\big{\|}}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}+{\big{\|}Bi(z)\big{\|}}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}\bigg{)}$
$\qquad\qquad\qquad\quad\leq C|\xi|^{-\frac{1}{6}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{2}}e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
${\big{\|}\mu_{M}(\xi,z)\big{\|}}_{L_{z}^{2}(y,M)}\leq C{\big{\|}[1+|\xi|^{\frac{1}{3}}z]^{-\frac{1}{4}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}z]^{\frac{3}{2}}}\big{\|}}_{L_{z}^{2}(y,M)}$
$\qquad\qquad\qquad\qquad\leq C|\xi|^{-\frac{1}{6}}{\big{\|}[1+\tilde{z}]^{-\frac{1}{4}}e^{-\frac{2}{3}\tilde{z}^{\frac{3}{2}}}\big{\|}}_{L_{\tilde{z}}^{2}(({\rm Re\ }[(i\xi)^{\frac{1}{3}}y],M))}$
$\qquad\qquad\qquad\qquad\leq C|\xi|^{-\frac{1}{6}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{2}}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}.$
Combining (3.3), (3.4), (3.8), (3.12) with (3.13), we obtain
(3.14) $|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
$\qquad\qquad\qquad\qquad\qquad\quad+\big{|}\mu_{0}(\xi,y)\big{|}\cdot\int_{y}^{M}\big{|}\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz\bigg{\}}$
$\qquad\leq C|\xi|^{-\frac{1}{3}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([0,M])}$
$\qquad\qquad\cdot\bigg{\{}e^{-\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\cdot{\big{\|}\mu_{0}(\xi,z)\big{\|}}_{L_{z}^{2}([0,y])}$
$\qquad\qquad\quad+e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}}\cdot{\big{\|}\mu_{M}(\xi,z)\big{\|}}_{L_{z}^{2}([y,M])}\bigg{\}}$
$\qquad\leq C|\xi|^{-\frac{1}{2}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{3}{4}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
Hence, we have
(3.15) ${\big{\|}(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big{\|}}_{L_{y}^{2}(0,M)}\leq C(\alpha^{\prime})|\xi|^{-\frac{1}{2}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}(0,M)}\cdot|\xi|^{-\frac{1}{6}}$
$\qquad\leq C|\xi|^{-\frac{2}{3}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}(0,M)}\leq CM^{2}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}(0,M)},$
for $\alpha^{\prime}\in[0,\frac{1}{4})$. By (3.11) and (3.15), we obtain
(3.16) ${\big{\|}f_{e}\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f_{e}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{2}{\big{\|}h_{e}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{2}{\big{\|}h_{e}\big{\|}}_{L^{2}(\Omega_{M})}.$
Here $C$ is independent of $M$ and $f_{e}$.
Step 3. We derive the high-frequency estimates of $f_{e}$. For any $\bar{\alpha}\geq 0$, we set
(3.17) $B_{0}^{2}=|\xi|^{2+\bar{\alpha}}{\big{\|}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big{\|}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}.$
Multiplying the equation (3.2) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we have
(3.18) $i|\xi|^{2+\bar{\alpha}}{\big{\|}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
Here we use a fact that $\hat{f}_{e}\big{|}_{y=M}=0$. Set
(3.19) $B_{1}^{2}=$ $\ |\xi|^{1+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2},$
$B_{2}^{2}=$ $\ |\xi|^{\frac{5}{3}+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}.$
Then we have
(3.20) $|\xi|^{2+\bar{\alpha}}{\big{\|}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+B_{1}^{2}\leq\big{|}-\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big{|}$
$\qquad\leq\big{|}-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big{|}$
$\qquad\leq C\bigg{(}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}\bigg{)}B_{0}$
$\qquad\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}).$
Here $C$ is independent of $\bar{\alpha}$. Multiplying the equation (3.2) by $|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}$, we obtain
(3.21) $i\xi|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}=\langle\hat{h}_{e},|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}\rangle,\quad\mbox{by}\ \ \hat{f}_{e}(\cdot,M)=0.$
Then we get
(3.22) $|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}\leq{\big{\|}|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}$
$\qquad\leq CB_{0}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}.$
Thus, by (3.19), we obtain
(3.23) $B_{2}^{2}\leq$ $\ CB_{0}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}.$
On the other hand, it is straight to obtain
(3.24) $|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big{\|}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}=|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}\partial_{y}(y)|\hat{f}_{e}|^{2}dy$
$\qquad=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big{|}_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
$\qquad\leq 2{\big{\|}|\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}$
$\qquad\leq CB_{0}B_{2}.$
By (3.17), (3.19), (3.20), (3.23), and (3.24), we have
(3.25) $B_{0}^{2}+B_{1}^{2}\leq C|\xi|^{\bar{\alpha}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}^{2}.$
Then by (3.23) and (3.25), there holds
(3.26) $B_{2}^{2}\leq$ $\ CB_{0}^{2}+C|\xi|^{\bar{\alpha}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}^{2}\leq C|\xi|^{\bar{\alpha}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}^{2}.$
Applying Plancherel Theorem, we obtain
(3.27) ${\big{\|}f_{e}\big{\|}}_{Y_{\bar{\alpha}}}^{2}\leq$ $\ C\bigg{(}{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}y\partial_{x}f_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f_{e}\big{\|}}_{L^{2}}^{2}$
$\quad+{\big{\|}|\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}f_{e}\big{\|}}_{L^{2}}^{2}\bigg{)}$
$=$ $\ C\bigg{(}{\big{\|}|\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\xi|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L^{2}}^{2}$
$\quad+{\big{\|}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L^{2}}^{2}+{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\hat{f}\big{\|}}_{L^{2}}^{2}\bigg{)}$
$\leq$ $\ C\bigg{(}\int_{{\mathbb{R}}}B_{0}^{2}+B_{1}^{2}+B_{2}^{2}+B_{0}^{2}+{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}^{2}d\xi\bigg{)}$
[…]
$=$ $\ C{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}h_{e}\big{\|}}_{L^{2}}^{2}.$
By (3.16) and (3.27), we get
(3.28) ${\big{\|}f_{e}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C\big{[}(1+M^{2}){\big{\|}h_{e}\big{\|}}_{L^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big{\|}}_{L^{2}}\big{]}.$
Here $C$ is independent of $\alpha,M$, and $f_{e}$.
Remark 3.1.
The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows.
(3.29) $Ai(z)=\frac{\sqrt{z}}{3}\big{(}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big{)},\quad z\in{\mathbb{C}},$
$Bi(z)=\sqrt{\frac{z}{3}}\big{(}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})+I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big{)},\quad z\in{\mathbb{C}},$
$I_{\nu}(z)=e^{-\frac{\nu\pi i}{2}}J_{\nu}(ze^{\frac{\pi i}{2}}),\quad-\pi\leq\text{Arg}(z)\leq\frac{\pi}{2},\ \nu\in{\mathbb{R}}.$
Here $J_{\nu}(z)=\sum\limits_{k=0}^{\infty}\frac{(-1)^{k}\big{(}\frac{z}{2}\big{)}^{2k+\nu}}{k!\Gamma(k+\nu+1)}$ denotes the Bessel function satisfying the Bessel equation
$\big{(}z^{2}\frac{d^{2}}{dz}+z\frac{d}{dz}+(z^{2}-\nu^{2})\big{)}J_{\nu}(z)=0.$
We then derive
(3.30) ${\mathcal{C}}(z)$= $\ \sqrt{3}Ai(z)+Bi(z)=\sqrt{\frac{4z}{3}}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})$
$=$ $\ \sqrt{\frac{4z}{3}}e^{\frac{\pi i}{6}}J_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}}e^{\frac{\pi i}{2}}),\qquad-\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}.$
For $\nu>-1$, the Bessel function $J_{\nu}(z)$ has only real zeros in $\mathbb{C}$} [31, Page 482]. Then by (3.30), the zeros of ${\mathcal{C}}(z)$ in the sector $\big{\{}z\in{\mathbb{C}}|\ -\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}\big{\}}$ lie only on rays $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=-\frac{\pi}{3}\ \text{or}\ \frac{\pi}{3}\big{\}}$. Note that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big{\}}$.
By homogenization of (1.15), it is straight to obtain the linearized problem:
(3.31) $\left\{\begin{aligned} &y\partial_{x}f-\partial_{y}^{2}f=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[f]=A+F,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
In order to solve the problem (3.31), we first omit the integral condition.
Lemma 3.2.
Let $\bar{\alpha}\geq 0$ and $g=\partial_{x}|\partial_{x}|A\in H^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}\cap\dot{H}^{-1}$. Then, there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ of the boundary-value problem:
(3.32) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big{|}_{y=0}=g,\quad x\in{\mathbb{R}},\end{aligned}\right$.
with
(3.33) ${\big{\|}\omega_{b}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big{\|}}_{L^{2}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})},$
where $C>0$ is a constant independent of $\bar{\alpha}$ and $A$.
Proof.
To solve the boundary-value problem (3.32), we first take the Fourier transform in variable $x$. Then we obtain
(3.34) $\left\{\begin{aligned} i\xi y\hat{\omega}_{b}-\partial_{y}^{2}\hat{\omega}_{b}=&\ 0,\quad(\xi,y)\in{\mathbb{R}}\times(0,M)\\ \partial_{y}\hat{\omega}_{b}\big{|}_{y=0}=\hat{g}=&\ i\xi|\xi|\hat{A},\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
The general solution to this BVP (3.34) is
(3.35) $\hat{\omega}_{b}(\xi,y)=C(\xi)\bigg{(}\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg{)}+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}\hat{g}$.
Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula as following:
(3.36) $\hat{\omega}_{b}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi<0.\end{aligned}\right$.
Then we have
(3.37) ${\big{\|}(|\partial_{x}|^{\frac{1}{3}}y)^{\alpha^{\prime}}\omega_{b}\big{\|}}_{L^{2}(\Omega_{M})}\leq C{\big{\|}(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{\omega}_{b}\big{\|}}_{L^{2}(\Omega_{M})}$
$\qquad\leq C{\big{\|}{\big{\|}Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big{\|}}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big{\|}}_{L_{\xi}^{2}}$
$\qquad\leq C{\big{\|}|\xi|^{\frac{3}{2}}\hat{A}(\xi)\big{\|}}_{L_{\xi}^{2}}=C{\big{\|}|\partial_{x}|^{\frac{3}{2}}A\big{\|}}_{L^{2}({\mathbb{R}})}$
$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}A\big{\|}}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
Since $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
(3.38) ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}=$ $\ {\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}$
$\ \ +{\big{\|}|\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big{\|}}_{L_{x,y}^{2}}$
$\leq$ $\ C{\big{\|}|\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big{\|}}_{L^{2}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{13+3\bar{\alpha}}{6}}A\big{\|}}_{L^{2}({\mathbb{R}})}.$
It follows $\omega_{b}\in X_{\bar{\alpha}}$ from (3.37) and (3.38). In particular, we have
(3.39) ${\big{\|}\omega_{b}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
Now, we are in the position to prove Theorem 1.1. Denote $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to
(3.40) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\\ &u_{0}=I_{y}[\omega_{0}],\quad v_{0}=-\partial_{x}I_{y}[u_{0}],\\ &\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0}(x),\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x).\end{aligned}\right$.
Unless otherwise specified, the parameters $\alpha,\bar{\alpha}$ satisfy $\alpha>0$ and $\bar{\alpha}\geq 0$.
Proof of Theorem 1.1.
We hope that we take (3.36) as a solution to (3.40), that is,
(3.41) $\hat{\omega}_{0}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}(i\xi|\xi|\hat{A}_{0})=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi<0.\end{aligned}\right$.
Here $\hat{f}(\xi,y)={\mathcal{F}}_{x\to\xi}[f](\xi,y)$. It is straight to verify that
(3.42) $\omega_{0}(\cdot,y)\big{|}_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A_{0}.$
Substituting the formula (3.41) into the integral condition $I_{M}[\hat{\omega}_{0}]=\hat{A}_{0}+\hat{F}$, we have
(3.43) $\hat{A}_{0}+\hat{F}$= $\ \int_{0}^{M}\hat{\omega}_{0}(\xi,y)dy$
$=$ $\ \int_{0}^{M}Ai(e^{\frac{{\rm sgn}\ \xi}{6}\pi i}|\xi|^{\frac{1}{3}}y)dy\cdot\frac{1}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{6}\pi i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}$
$=$ $\ \int_{0}^{e^{\frac{{\rm sgn}\ \xi}{6}\pi i}|\xi|^{\frac{1}{3}}M}Ai(z)dz\cdot\frac{1}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\hat{A}_{0}$
$=$ $\ \frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn}\ \xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\hat{A}_{0}.$
Here
(3.44) $p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
Indeed, the integral is well-defined since $Ai(z)$ is analytic. We set
(3.45) $m(\xi)=$ $1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}.$
Then by (3.43), we obtain
(3.46) $\hat{A}_{0}=-\frac{1}{m(\xi)}\hat{F}$.
It follows from (3.41) that
(3.47) $\hat{\omega}_{0}(\xi,y)=-\frac{Ai(e^{\frac{{\rm sgn\ }\xi}{6}\pi i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{6}\pi i}}\frac{i\xi|\xi|^{\frac{2}{3}}}{m(\xi)}\hat{F}(\xi).$
To ensure that the formulas (3.46) and (3.47) are well-defined, we need $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. Then, we give an estimate of $m(\xi)$. We define
$p_{0}:=\mathop{sup}\limits_{{\rm Arg\ }z\in\{\pm\frac{\pi}{6}\}}\bigg{|}\int_{0}^{z}Ai(\zeta)d\zeta\bigg{|}.$
Note that $0\leq p_{0}<\infty$, then we obtain
$a_{0}:=\bigg{(}\frac{4p_{0}}{3|Ai^{\prime}(0)|}\bigg{)}^{-\frac{3}{4}}>0.$
It follows that
(3.48) $m(\xi)\geq\frac{1}{4},\quad\forall\ |\xi|\leq a_{0}.$
Using the following identity
(3.49) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3}.$
As $M\to\infty$, the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$. By calculations, we have
(3.50) $\bigg{|}1-\frac{1}{3Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg{|}^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
Then there exists a large enough constant $M_{0}>0$, when $M\geq M_{0}$, by (3.48) and (3.50), we also obtain
(3.51) $|m(\xi)|^{-1}=\bigg{|}1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg{|}^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
It concludes that $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, by Fourier multiplier theorem, there exists a constant $C>0$ such that
(3.52) ${\big{\|}m(D)^{-1}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}f\big{\|}}_{L^{2}({\mathbb{R}})},\quad\forall\ f\in L^{2}({\mathbb{R}}).$
Using Lemma 3.2 and the estimate (3.52), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as following:
(3.53) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big{\|}}_{L^{2}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}F\big{\|}}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we obtain
(3.54) ${\big{\|}u_{0}\big{\|}}_{L^{\infty}(\Omega_{M})}+{\big{\|}A_{0}\big{\|}}_{L^{\infty}({\mathbb{R}})}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big{\|}\omega_{0}\big{\|}}_{X_{0}}.$
It completes the proof. ∎
Remark 3.2.
Note that ${\rm Re\ }\bigg{(}\frac{p_{M}(\xi)}{e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}\bigg{)}>0$ for $|\xi|^{\frac{1}{3}}M>0$ by numberical results (see figure 2, where $r=|\xi|^{\frac{1}{3}}M$). Then by continuity of $m(\xi)$, we have $\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}$ for any $\xi\in{\mathbb{R}}$ and $M>0$, where $C>0$ depends on $M^{-1}$.
Figure 2. Analysis about $m(\xi)$.
4. Proof of Main Theorem
Let us recall the perturbation system (1.15) that
(4.1) $\left\{\begin{aligned} &y\omega_{x}-\omega_{yy}=-u\omega_{x}-v\omega_{y},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\omega_{y}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
In order to find a solution of the equation (4.1), we use the iteration method. Set
$(\bar{\omega},\bar{A},\bar{u},\bar{v})=(\omega-\omega_{0},A-A_{0},u-u_{0},v-v_{0}).$
Then, we reduce (4.1) to the following BVP:
(4.2) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(\bar{u}+u_{0})\partial_{x}(\bar{\omega}+\omega_{0})-(\bar{v}+v_{0})\partial_{y}(\bar{\omega}+\omega_{0}),\\ &\bar{u}=I_{y}[\bar{\omega}],\quad\bar{v}=-\partial_{x}I_{y}[\bar{u}],\\ &\partial_{y}\bar{\omega}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A}.\end{aligned}\right$.
Thus, we can consider the iterated system as following:
(4.3) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{k+1}-\partial_{y}^{2}\bar{\omega}_{k+1}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}),\\ &\bar{u}_{k+1}=I_{y}[\bar{\omega}_{k+1}],\quad\bar{v}_{k+1}=-\partial_{x}I_{y}[\bar{u}_{k+1}],\\ &\bar{u}_{k}=I_{y}[\bar{\omega}_{k}],\quad\bar{v}_{k}=-\partial_{x}I_{y}[\bar{u}_{k}],\\ &\partial_{y}\bar{\omega}_{k+1}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\quad I_{M}[\bar{\omega}_{k+1}]=\bar{A}_{k+1}.\end{aligned}\right$.
Recall that the space $X_{\alpha,\infty}$ is defined as
(4.4) $X_{\alpha,\infty}=\bigg{\{}f:{\mathbb{R}}\to{\mathbb{R}}\big{|}\ \ {\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{L^{2}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
First of all, we give some results about the iterated system (4.3).
Proposition 4.1.
Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$. Suppose $k$ is a non-negative integer and $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ of (4.3). Moreover, for $\varepsilon\in(0,\frac{1}{6}]$, we have
(4.5) ${\big{\|}\bar{\omega}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\big{[}\big{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\big{)}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\big{]}.$
Here $C=C(\alpha,\varepsilon)>0$ is a constant.
Proof.
Set $\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0})$, we decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
(4.6) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big{|}_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big{|}_{y=M}=0,\end{aligned}\right$.
$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1}.\end{aligned}\right$.
with
(4.7) $I_{M}[\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}]=\bar{A}_{k+1}.$
By Lemma 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the following form
(4.8) $\bar{\omega}_{e,k+1}=$ $\ \pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg{\{}\mu_{M}(\xi,y)\int_{0}^{y}\mu_{0}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz$
$\qquad\qquad\qquad+\mu_{0}(\xi,y)\int_{y}^{M}\mu_{M}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz\bigg{\}}.$
By Lemma 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ as following:
(4.9) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
In view of (4.7), we have
(4.10) $\bar{A}_{k+1}=m(D)^{-1}I_{M}[\bar{\omega}_{e,k+1}].$
Here $m=m(\xi)$ is defined in (3.45).
By Lemma 3.1, Theorem 1.1, and Proposition 2.3, we obtain
(4.11) ${\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq$ $\ C\big{(}(1+M^{2}){\big{\|}\bar{h}_{k}\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\bar{h}_{k}\big{\|}}_{L^{2}(\Omega_{M})}\big{)}$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{0}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\partial_{x}A_{0}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
$\qquad+{\big{\|}\partial_{x}A_{k}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big{\|}\partial_{x}F\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2})$
$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}).$
By Lemma 3.2 and Proposition 2.4, we have
(4.12) ${\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}={\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
[…]
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}},$
${\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Combining (4.11) with (4.12), one gets
(4.13) ${\big{\|}\bar{\omega}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}$
$\qquad\leq{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\bigg{[}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2})+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\bigg{]}.$
Then it completes the proof.
To get the uniform estimates of the solutions, we use the following results which can be got the induction method.
Lemma 4.1.
Let $\alpha_{i},\beta_{i}\ (i=1,2)$ be positive constants. Assume that $\{a_{k}\}_{k\geq 0},\ \{b_{k}\}_{k\geq 1}$ are two non-negative sequences satisfying
(4.14) $(1)\ a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2},\quad\mbox{for any }k\geq 0;\quad(2)\ b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2},\quad\mbox{for any }k\geq 1.$
Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$, then we have
(4.15) $0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
Now, we are in the position to give the proof of Theorem 1.2.
Proof of Theorem 1.2.
The proof proceeds in three steps.
Step 1. Bounded sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
Taking $\bar{\omega}_{0}=\bar{A}_{0}=0$, by Proposition 4.1, we obtain a sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$. Set $a_{k}={\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}$. Clearly, we have $a_{0}=0$. By Proposition 4.1, one gets
(4.16) $a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
Using Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
(4.17) $\sup_{k\geq 0}\bigg{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
Step 2. Convergence of this sequence.
Set
[…]
By subtracting the equation (4.3) corresponding to $\bar{\omega}_{k+1}$ from the equation (4.3) corresponding to $\bar{\omega}_{k}$, we obtain
[…]
We set
[…]
In view of Remark 2.1, we have
[…]
It follows from (4.19) and (4.20) that
[…]
By Lemma 4.1, if $CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2}$ and
[…]
which means ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$. Then there holds
[…]
Step 3. Uniqueness of $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
Recall that
${\mathcal{X}}_{\alpha,M}$= $\ \bigg{\{}(\omega,A)\big{|}\ {\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}={\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A\big{\|}}_{X_{\alpha,\infty}}<\infty,$
$\qquad\qquad\quad\omega(\cdot,y)\big{|}_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg{\}}.$
By Theorem 1.1, we obtain $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$. Combing (4.8)-(4.9) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we have $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
(4.22) $\hat{\bar{\omega}}(\xi,y)\big{|}_{y=M}=\hat{\bar{\omega}}_{b}(\xi,y)\big{|}_{y=M}$
$\qquad=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0,\end{aligned}\right$.
which implies $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$. Moreover, it is evident to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}-\{0\}$ due to $F\not\equiv 0$, which implies the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial. Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions of the perturbation system (1.15). Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, then there holds
[…]
Similar to the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
(4.24) $\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big{|}_{y=0}=0,\quad\tilde{\omega}_{e}\big{|}_{y=M}=0,\end{aligned}\right$.
[…]
By Lemma 3.1, we have
(4.25) ${\big{\|}\tilde{\omega}_{e}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq$ $\ CM^{2}{\big{\|}\tilde{h}\big{\|}}_{L^{2}(\Omega_{M})}+C{\big{\|}\tilde{h}\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}.$
By Lemma 3.2, we obtain
(4.26) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\bigg{(}\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg{)}+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
In view of $\hat{\tilde{\omega}}_{e}\big{|}_{y=M}=0$, one gets
(4.27) $\hat{\tilde{\omega}}_{b}\big{|}_{y=M}=\hat{\tilde{\omega}}\big{|}_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$.
Combining (4.26) with (4.27), we obtain
$C(\xi)=0.$
Thus, we have
(4.28) $\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
This formula is identical to (3.36), then using the Lemma 3.2, we obtain
(4.29) ${\big{\|}\tilde{\omega}_{b}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}\tilde{A}\big{\|}}_{H^{\frac{4}{3}+\frac{\alpha}{2}}}\leq C{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}.$
Since $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
(4.30) $\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
It follows from Proposition 2.4 that
(4.31) ${\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\leq C{\big{\|}\tilde{\omega}_{e}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
By Remark 2.1, we obtain
(4.32) ${\big{\|}\tilde{h}\big{\|}}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}},$
${\big{\|}\tilde{h}\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
$\qquad\qquad\quad+CM\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
On the other hand, we know
(4.33) ${\big{\|}\tilde{A}\big{\|}}_{L^{\infty}({\mathbb{R}})}\leq{\big{\|}\tilde{\omega}\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
Combining (4.25), (4.29), (4.31), (4.32) with (4.33), we obtain
(4.34) ${\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\qquad\cdot\big{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}$
$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
$\qquad\qquad\cdot\big{(}{\big{\|}(\omega_{1},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}(\omega_{2},A_{2})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}.$
This implies the uniqueness of solution $(\omega,A)$ in ${\mathcal{X}}_{\alpha,M}$ to (1.15) when ${\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{\delta_{0}}{M^{\frac{5+3\varepsilon}{2}}}$ with $\delta_{0}=\delta_{0}(\alpha,\varepsilon)\ll 1$.
Appendix
This appendix provides a brief overview of the triple-deck formalism, based on the physical model of an incompressible boundary layer with a roughness element; for further details, see [26]. The roughness is located at a distance $L$ downstream from the leading edge. Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively. For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$. The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge. The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and its streamwise length scale is $O(\epsilon^{3})$. Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (near $X=1$). The three distinct layers are summarized as follows:
I. Viscous sublayer: Vertical scale $O(\epsilon^{5})$, streamwise scale $O(\epsilon^{3})$.
II. Main layer: Vertical scale $O(\epsilon^{4})$, streamwise scale $O(\epsilon^{3})$.
III. Upper layer: Vertical scale $O(\epsilon^{3})$, streamwise scale $O(\epsilon^{3})$.
These scales arise from matched asymptotic expansions, connecting the triple-deck theory to classical boundary layer theory in the outer flow region. For clarity, the spatial arrangement of layers is illustrated in Figure 3, showing the viscous sublayer I adjacent to the wall and the upper layer III in the potential flow region.
Figure 3. Demonstration of the triple-deck structure (not to scale): I Viscous Sublayer; II Main Layer; III Upper Layer; $\delta$ Thickness of Prandtl layer; $\tilde{U}_{0}$ Streamwise velocity.
It is well-known that the 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
(A.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
In the main layer (layer II), we introduce a pair of local coordinates,
$x_{*}=\epsilon^{-3}(X-1)=O(1),\quad\tilde{y}=\epsilon^{-4}Y=O(1).$
The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+O(\epsilon^{4})$ and a mean-flow distortion,
(A.2) $u_{NS}(X,Y;\epsilon)=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon\tilde{u}(x_{*},\tilde{y})+o(\epsilon),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\tilde{v}(x_{*},\tilde{y})+o(\epsilon^{2}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\tilde{p}(x_{*},\tilde{y})+o(\epsilon^{2}).$
Substituting the expansions (A.2) into (A.1) and collecting the leading-order terms, we derive the following system
(A.3) $\left\{\begin{aligned} \tilde{U}_{0}\partial_{x_{*}}\tilde{u}+\tilde{U}_{0}^{\prime}\tilde{v}=&\ 0,\\ \partial_{\tilde{y}}\tilde{p}=&\ 0,\\ \partial_{x_{*}}\tilde{u}+\partial_{\tilde{y}}\tilde{v}=&\ 0.\end{aligned}\right$.
The solutions of this system are
(A.4) $\tilde{u}=\lambda^{-1}A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad\tilde{v}=-\lambda^{-1}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad\tilde{p}=\tilde{p}(x_{*}).$
Here $\lambda:=\tilde{U}_{0}^{\prime}(0)>0$ and $A_{*}$ is an unknown function (typically referred to as a displacement of velocity) satisfying $A_{*}(-\infty)=0$.
As $\tilde{y}\to\infty$, we find that $\tilde{v}$ approaches $\lambda^{-1}A_{*}^{\prime}(x_{*})$, which does not satisfy the far-field attenuation condition, $\lim\limits_{Y\uparrow\infty}v_{NS}(X,Y)=0$. This indicates the emergence of an upper layer (layer III in Figure 3). In this upper layer, we introduce the local coordinate
$\bar{y}=\epsilon^{-3}Y=O(1).$
Matching with the upper limit of the main-layer solutions, we can express the flow field as
(A.5) $u_{NS}(X,Y;\epsilon)=$ $1+\epsilon^{2}\bar{u}(x_{*},\bar{y})+o(\epsilon^{2}),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\bar{v}(x_{*},\bar{y})+o(\epsilon^{2}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}\bar{p}(x_{*},\bar{y})+o(\epsilon^{2}).$
Substituting the expansions (A.5) into (A.1) and retaining the leading-order terms, we derive
[…]
The far-field attenuation condition indicates
(A.7) $\lim_{\sqrt{x_{*}^{2}+\bar{y}^{2}}\to\infty}(\bar{u},\bar{v},\bar{p})=\mathbf{0}.$
Matching with the main layer, we derive
(A.8) $\bar{v}(x_{*},0)=\lim_{\tilde{y}\uparrow\infty}\tilde{v}=-\lambda^{-1}A_{*}^{\prime}(x_{*}),\quad\bar{p}(x_{*},0)=\tilde{p}(x_{*}).$
Since $(\bar{p},\bar{v})$ satisfies the Cauchy-Riemann equations (cf. (A.6)), let us define $f(z)=\bar{p}(x_{*}\bar{y})+i\bar{v}(x_{*},\bar{y})$ with $z=x_{*}+i\bar{y}$. Then, we obtain
(A.9) $f(z)=$ $\ \frac{1}{2\pi i}\int_{{\mathbb{R}}}\frac{f(\zeta)-\overline{f(\zeta)}}{\zeta-z}d\zeta=\ -\frac{1}{\pi\lambda}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-z}d\zeta,\quad\forall\ \text{Im }z>0.$
This implies that
(A.10) $\tilde{p}(x_{*})=\bar{p}(x_{*},0)=\lim_{\text{Im}(z)\to 0^{+}}\text{Re}\ f(z)=-\frac{1}{\pi\lambda}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
Now let us consider the lower boundary of the main-layer solution. As $\tilde{y}\to 0^{+}$, we find that $\tilde{u}$ approaches $A_{*}(x_{*})$, disagree with the no-slip condition. Thus, we need to consider the viscous sublayer (layer I in Figure 3). For convenience, we introduce
$y_{*}=\epsilon^{-5}Y=O(1).$
Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$. Thus, the main-layer solution behaves like $u_{NS}=\bar{U}_{0}+\epsilon\tilde{u}+o(\epsilon)\to\epsilon(\lambda y_{*}+A_{*})+o(\epsilon)$, which determines the magnitude of the sublayer-layer solution. From the balance of the NS equations, we can expand the solution as,
(A.11) $u_{NS}(X,Y;\epsilon)=$ $\ \epsilon u_{*}(x_{*},y_{*})+o(\epsilon),$
$v_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{3}v_{*}(x_{*},y_{*})+o(\epsilon^{3}),$
$p_{NS}(X,Y;\epsilon)=$ $\ \epsilon^{2}P_{*}(x_{*})+o(\epsilon^{2}).$
Here, we have imposed the fact that, to leading order, the pressure remains uniform across the thin sublayer. Substituting (A.11) into (A.1) and collecting the leading-order terms, we derive the following system:
(A.12) $\left\{\begin{aligned} u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=&\ 0,\\ \partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=&\ 0.\end{aligned}\right$.
The velocity field is subject to the no-slip, non-penetration conditions,
(A.13) $u_{*},v_{*}]\big{|}_{y_{*}=F(x_{*})}=0.$
In the upstream limit, the velocity approaches the unperturbed Blasius profile,
(A.14) $(u_{*}-\lambda y_{*})\big{|}_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big{|}_{x_{*}\downarrow-\infty}=0.$
The upper boundary condition of the sublayer equations should be obtained by matching with the main layer. Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$. Then for any $\beta_{0}>0$, we have
(A.15) $u_{NS}=$ $\ \lambda\tilde{y}+\epsilon\tilde{u}(x_{*},0)+o(\epsilon+\tilde{y}^{4-\beta_{0}})=\lambda\tilde{y}+\epsilon A_{*}(x_{*})+o(\epsilon+\tilde{y}^{4-\beta_{0}})$
$=$ $\ \epsilon\big{(}\lambda y_{*}+A_{*}(x_{*})\big{)}+o\big{(}\epsilon+\epsilon^{(4-\beta_{0})(1-\alpha_{0})}\big{)}.$
To ensure the second part to be negligible, we require $(4-\beta_{0})(1-\alpha_{0})\geq 1$, namely, $\alpha_{0}\in(0,3/4)$ and $\beta_{0}\in(0,4)$. Likewise, we can derive that the leading-order pressure in the sublayer and that in the main layer are equal. Subsequently, we obtain
(A.16) $\big{(}u_{*}(x_{*},y_{*})-\lambda y_{*}\big{)}\big{|}_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad\tilde{p}(x_{*})=P_{*}(x_{*}).$
From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as following:
(A.17) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &[u_{*},v_{*}]\big{|}_{y_{*}=F(x_{*})}=0,\quad(u_{*}-\lambda y_{*})\big{|}_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad x\in{\mathbb{R}},\\ &(u_{*}-\lambda y_{*})\big{|}_{x_{*}\downarrow-\infty}=0,\quad y_{*}\in{\mathbb{R}}_{+},\\ &P_{*}(x_{*})=-\frac{1}{\pi\lambda}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
Here $\Omega_{\infty,F}=\big{\{}(x_{*},y_{*})\big{|}\ x_{*}\in{\mathbb{R}},\ y_{*}>F(x_{*})\big{\}}$ and $\lim_{x_{*}\downarrow-\infty}F(x_{*})=0$.
Indeed, the above nonlinear system is usually solved using numerical approach. In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$. Consequently, the solution $\phi_{*}=(u_{*},v_{*},P_{*})$ should be denoted by $\phi_{*,\epsilon}$ to include the impact of the upper boundary. It is clear that $\phi_{*,\epsilon}$ also satisfy the system (A.17), except for the upper boundary condition (A.16), which is converted to
(A.18) $(u_{*,\epsilon}(x_{*},y_{*})-\lambda y_{*})\big{|}_{y_{*}=f(x_{*})}=A_{*}(x_{*}).$
When there is no ambiguity, we still denote $\phi_{*,\epsilon}$ as $\phi_{*}$ for simplicity.
There is a useful operation named the Prandtl transformation for solving the system (A.17), which takes the following form:
(A.19) $y=y_{*}-F(x_{*}),\ v_{*}=V+u_{*}F^{\prime}(x_{*}).$
This transformation merely shifts down the domain by a distance $F(x_{*})$ in the $y$-direction, with the governing equations being unchanged. It is preferable to take $f(x_{*})=\varepsilon^{-\alpha_{0}}+F(x_{*})$, then, the boundary conditions in the transverse direction are changed to
(A.20) $(u_{*},V)\Big{|}_{y=0}=0,\quad(u_{*}-\lambda y)\Big{|}_{y=\epsilon^{-\alpha_{0}}}=A_{*}+\lambda F.$
Typically, we take $\lambda=1$ in this paper.
Acknowledgments. M. Dong is partially supported by the Chinese Academy of Sciences (CAS) Project for Young Scientists in Basic Research under Grant No. YSBR-087. C. Wang is partially supported by the National Natural Science Foundation of China (NSFC) under Grant No. 12471189. Z. Zhang is partially supported by the NSFC under Grant No. 12288101.

### v2

@@
-Existence and Uniqueness of Solutions for Steady Triple-Deck Equations
+On the classical solution for the steady triple-deck equations
 Abstract.
-Triple-deck structures arise in boundary-layer flows when small localized wall perturbations induce moderate mean-flow distortion.
-Most existing studies in this field focus on the derivations of the asymptotic and numerical solutions of this system, while rigorous mathematical results remain scarce.
-In this paper, we establish, for the first time, the existence and uniqueness of the solutions to the steady triple-deck equations.
-The steady triple-deck equations can be reduced to an Airy equation with frequency as a parameter via Fourier transformation.
-The main difficulty is that the Airy equation degenerates at zero frequency, hindering uniform estimates of the solution with respect to the frequency parameter.
-A key contribution is that we find a subtle function involving the distribution of zeros of the Bessel function to overcome the degeneracy by analyzing the structure of the Airy function.
-Additionally, we derive a higher-frequency estimate and modify an elliptic multiplier to obtain uniform estimates of the solution.
+This paper establishes the existence and uniqueness of classical solutions to the steady Triple-Deck equations, which describe incompressible boundary layer flow over localized roughness at high Reynolds numbers.
+The triple-deck theory was developed to overcome the Goldstein singularity in classical Prandtl boundary layer theory, capturing the interaction between the viscous sublayer, main layer, and upper layer when surface roughness of height $O({\rm Re}^{-\frac{5}{8}})$ is present.
+The key ingredients of this paper include: (1) A decomposition separating roughness effects from nonlinear difficulties; (2) A novel Green’s function using Airy functions that overcomes low-frequency singularities via the non-vanishing property of $\sqrt{3}Ai(z)+Bi(z)$; (3) The introduction of weighted Sobolev norms ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$ of the vorticity yielding $M$-independent estimates for displacement $A$.
+As a byproduct, local uniqueness of Couette flow is established when $F=0$.
 1. Introduction
+1.1. The Triple-Deck equations
 In 1904, Prandtl [21] proposed the boundary layer theory to address attached flow problems, achieving remarkable success in fluid dynamics research.
@@
 Notably, Landau and Lifshitz [14] revealed that the velocity component normal to the body surface grows unboundedly (termed the Goldstein singularity), which hinders the solution from being further extended to the reverse-flow region [6].
-To overcome these difficulties, scholars including A.F.
-Messiter [17] and K.
-Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s.
+To overcome these difficulties, scholars including Messiter [17] and Stewartson [27] developed the triple-deck theory from the late 1960s to the early 1970s.
 In this theory, three distinct layers—a viscous sublayer, a main layer, and an upper layer—appear in the transverse direction, for which uniformly valid approximate solutions are derived by the asymptotic matched method.
-For instance, Messiter successfully eliminated the Goldstein singularity in his study on the flow near the trailing edge of an incompressible flat plate, demonstrating the theory’s mathematical resolution to singular behaviors [17].
-In physics, the triple-deck structure is observed when localized surface perturbations exist.
-For demonstration, we focus on the steady triple-deck flow over an incompressible two-dimensional (2D) flat plate with a localized roughness element.
-The roughness is located at a distance $L$ downstream from the leading edge.
-Taking $L$ as the reference length, the Reynolds number is defined as $\text{Re}=\frac{U_{\infty}L}{\nu}$, with $U_{\infty}$ and $\nu$ being the velocity of the incoming flow and the kinematic viscosity, respectively.
-For convenience, we introduce a small parameter $\epsilon=\text{Re}^{-\frac{1}{8}}\ll 1$.
-The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge.
-The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and we denote it as $\{(X,Y)\big{|}\ Y=\epsilon^{5}F(\frac{X-1}{\epsilon^{3}}),\ X\in{\mathbb{R}}\}$ for some given function $F=O(1)$ as $\epsilon\to 0$.
-Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (see figure 1).
+The 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
+(1.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
+Here $(u_{NS},v_{NS})$ denotes the velocity vector, $p_{NS}$ pressure, and $\epsilon={\rm Re}^{\frac{1}{8}}$ with the Reynolds number ${\rm Re}\gg 1$.
+In physics, the triple-deck structure emerges when localized surface roughness of height ${\mathcal{O}}(\epsilon^{5})$ is present.
+We then denote it as
+$\Big\{(X,Y)\big|\ Y=\epsilon^{5}F\big(\frac{X-1}{\epsilon^{3}}\big),\ X\in{\mathbb{R}}\Big\}$
+for some given function $F={\mathcal{O}}(1)$ as $\epsilon\to 0$.
+Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (see figure 1).
 Figure 1. Demonstration of the triple-deck structure (not to scale): I Viscous Sublayer; II Main Layer; III Upper Layer; $\delta$ Thickness of Prandtl layer.
-We denote the velocity vector by ${\mathbf{u}}_{NS}=(u_{NS}(X,Y),v_{NS}(X,Y))$, the pressure $p_{NS}$, and demonstrate their scales in Table 1.
+We demonstrate the scales of $(u_{NS},v_{NS},p_{NS})$ in Table 1.
 I II III
-$u_{NS}$ $O(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+O(\epsilon)$ $1+O(\epsilon^{2})$
-$v_{NS}$ $O(\epsilon^{3})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
-$p_{NS}$ $O(\epsilon^{2})$ $O(\epsilon^{2})$ $O(\epsilon^{2})$
+$u_{NS}$ ${\mathcal{O}}(\epsilon)$ $\tilde{U}_{0}(\frac{Y}{\epsilon^{4}})+{\mathcal{O}}(\epsilon)$ $1+{\mathcal{O}}(\epsilon^{2})$
+$v_{NS}$ ${\mathcal{O}}(\epsilon^{3})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
+$p_{NS}$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$ ${\mathcal{O}}(\epsilon^{2})$
 Table 1. Scales of velocities and pressure in the triple-deck structure: $\tilde{U}_{0}$ Streamwise velocity in Prandtl layer.
-It is well-known that the 2D steady incompressible Navier-Stokes equations take the following dimensionless form:
-(1.1) $\left\{\begin{aligned} &{\mathbf{u}}_{NS}\cdot\nabla{\mathbf{u}}_{NS}-\epsilon^{8}\Delta{\mathbf{u}}_{NS}+\nabla p_{NS}=0,\\ &\nabla\cdot{\mathbf{u}}_{NS}=0.\end{aligned}\right$.
-Based on the scaling relations in Figure 1 and Table 1, and by applying the method of matched asymptotic expansions (see details in Appendix), we derive that the flows in the upper layer and main layer are given by
-(1.2) $u_{NS}\sim 1-\epsilon\Re(f(z)),\quad v_{NS}\sim\epsilon^{2}\Im(f(z)),\quad p_{NS}\sim\epsilon^{2}\Re(f(z)),\quad z=x_{*}+i\bar{y},$
-$u_{NS}\sim\tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad v_{NS}\sim-\epsilon^{2}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad p_{NS}\sim\epsilon^{2}\tilde{p}(x_{*}).$
-Here $x_{*}=\frac{X-1}{\epsilon^{3}},\ \bar{y}=\frac{Y}{\epsilon^{3}},\ \tilde{y}=\frac{Y}{\epsilon^{4}}$, $\tilde{U}_{0}^{\prime}(0)=1$, $A_{*}$ is an unknown function (called the displacement of the velocity), and
-(1.3) $f(z)=-\frac{1}{\pi}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}}{\zeta-z}d\zeta,\quad\tilde{p}(x_{*})=\lim_{\Im(z)\to 0^{+}}\Re(f(z))=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
-For the viscous sublayer (region I), we derive the governing equations as
-(1.4) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\end{aligned}\right$.
+Based on the scaling relations in Figure 1 and Table 1, and via the method of matched asymptotic expansions applied to (1.1) (see Appendix for details), we derive the governing equations for the viscous sublayer (region I) as
+(1.2) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0.\end{aligned}\right$.
 Here $y_{*}=\frac{Y}{\epsilon^{5}}$ and $(\epsilon u_{*},\epsilon^{3}v_{*},\epsilon^{2}P_{*})$ are the leading-order terms of $(u_{NS},v_{NS},p_{NS})$.
@@
 (1.4) $(u_{*}-y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big|_{x_{*}\downarrow-\infty}=0.$
-To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution.
-Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$.
-Since the Blasius solution satisfies $\tilde{U}_{0}\to\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$, then for any $\beta_{0}>0$, it follows that
-(1.7) $u_{NS}=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y})+o(\epsilon)=\tilde{y}+\epsilon A_{*}(x_{*})+o(\epsilon+\tilde{y}^{4-\beta_{0}})$
-$=$ $\ \epsilon\big{(}y_{*}+A_{*}(x_{*})\big{)}+o\big{(}\epsilon+\epsilon^{(4-\beta_{0})(1-\alpha_{0})}\big{)}.$
-To ensure the second part to be negligible, we require $(4-\beta_{0})(1-\alpha_{0})\geq 1$, namely, $\alpha_{0}\in(0,\frac{3}{4})$ and $\beta_{0}\in(0,4)$.
-Likewise, we can derive that the leading-order pressure in the sublayer and that in the main layer are equal, that is,
+To close the system, the upper boundary condition of the sublayer equations is enforced by matching with the main layer solution (see Appendix for details).
+Then we can derive the leading-order pressure as follows
 (1.5) $P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.$
-In light of (1.7), it is permissible to set the upper boundary of the viscous sublyaer domain to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$, where $f$ is to be determined.
-Consequently, the upper boundary condition is
+And the upper boundary condition of $u_{*}$ is
 (1.6) $(u_{*}(x_{*},y_{*})-y_{*})\big|_{y_{*}=f(x_{*})}=A_{*}(x_{*}).$
-By (1.4)-(1.6), and (1.9), we derive the closed system for the viscous sublayer as
+Here $f$ is given (to be determined later). Consequently, by (1.2)-(1.6), we derive the closed system for the viscous sublayer as
 (1.7) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad F(x_{*})<y_{*}<f(x_{*}),x_{*}\in{\mathbb{R}},\\ &(u_{*},v_{*})\big|_{y_{*}=F(x_{*})}=0,\quad(u_{*}(x_{*},y_{*})-y_{*})\big|_{y_{*}=f(x_{*})}=A_{*}(x_{*}),\quad x_{*}\in{\mathbb{R}},\\ &(u_{*}-y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad P_{*}\big|_{x_{*}\downarrow-\infty}=0,\quad F(-\infty)<y_{*}<f(-\infty),\\ &P_{*}=-\frac{1}{\pi}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta.\end{aligned}\right$.
-A critical observation is that the flows (1.2) in the upper two layers only depend on the unknown quantities $(u_{*},v_{*},A_{*},P_{*})$ within the viscous sublayer, this paper thus focuses on the equations (1.10) of the viscous sublayer (henceforth referred to as the triple-deck equations).
-1.1. Known results
-Mathematically, the triple-deck equations share a structural resemblance to the Prandtl equations of classical boundary layers, yet they differ fundamentally in domain and pressure determination.
+The system (1.7) is referred to as the Triple-Deck equations.
+Recall that the Prandtl equations of classical boundary layers take
+(1.8) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad\text{in }(0,L_{0})\times{\mathbb{R}}_{+},\\ &(u_{*},v_{*})\big|_{y_{*}=0}=0,\quad x_{*}\in{\mathbb{R}}_{+},\\ &u_{*}\big|_{x_{*}=0}=u_{0},\quad y_{*}\in{\mathbb{R}}_{+},\end{aligned}\right$.
+where $u_{0}$ is prescribed.
+Mathematically, the Triple-Deck equations share a structural resemblance to the Prandtl equations (1.8), yet they differ fundamentally in domain and pressure determination.
 The Prandtl equations are defined over $(0,L_{0})\times\mathbb{R}_{+}$, with pressure $P_{*}$ dictated by external potential flow $u_{E}(x_{*})$ via the Bernoulli relation
 (1.9) $P_{*}+\frac{1}{2}u_{E}^{2}=\text{Const,}$
-which means that the pressure is a given function.
-There are a lot of results on this topic.
-The global existence of solutions to the Prandtl equations was first established by Oleinik [19], and subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30].
-Moreover, the flow separation and reverse phenomena have also attracted the attention of many scholars [2, 10, 15, 25].
-Compared with (1.11), the pressure of triple-deck model is given by (1.8) which is defined by unknown function $A_{*}$.
-Due to that, to construct a solution of triple-deck model is not an easy way.
-For the unsteady case, Iyer and Vicol [11], Gérard-Varet et al.
-[5] proved the well-posedness of solutions in real analytic spaces, while Dietert and Gérard-Varet [3] revealed instability for some smooth initial data, indicating that the class of analytic functions is optimal for well-posedness.
-For steady cases, mathematical results are sparser.
-When $F=0$, Iyer and Maekawa [9] proved a local Liouville-type theorem, pointing out that under specific rigid conditions, the equations admit only the Couette solution.
-However, physicists found that the triple-deck structure only appears in flow with some perturbations which means that $F\not\equiv 0$.
-For more results on physical analysis and numerical simulations, see [1, 4, 12, 26, 32].
-It is noteworthy that, for the triple-deck equations (1.10) in the presence of disturbance, the well-posedness of solutions lacks rigorous mathematical conclusions.
-This paper investigates the triple-deck equations (1.10) for some $f=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,\frac{3}{4})$, establishes the existence and uniqueness theory in the presence of disturbance for the first time, and verifies the compatibility between the disturbance scale and physical observation results.
+which means that the pressure is a given function. While, the pressure of Triple-Deck model is given by (1.5) with the unknown function $A_{*}$.
+The existence and uniqueness of classical solutions to the Prandtl equations were first established by Oleinik [19]; subsequent decades have yielded rich results on well-posedness, asymptotic behavior, and regularity [7, 8, 16, 23, 24, 29, 30].
+Moreover, flow separation and reverse-flow phenomena have attracted considerable attention from many scholars [2, 10, 15, 25].
+When $F=0$, it is easy to see that $(u_{*},v_{*},P_{*},A_{*})=(y,0,0,0)$ is a trivial solution to (1.7).
+In a recent work [9], Iyer and Maekawa proved the rigidity of this trivial solution under certain assumptions on the solution.
+However, physicists found that the triple-deck structure only appears when $F\not\equiv 0$.
+Since the $A_{*}$ remains unknown in this case, finding a non-trivial solution becomes a highly non-trivial problem.
+Numerous results are available on physical analysis and numerical simulations; see [1, 4, 12, 26, 32].
+Yet, if $F\not\equiv 0$, the well-posedness of (1.7) still lacks rigorous mathematical justification.
+For the unsteady case, Iyer and Vicol [11] and Gérard-Varet et al.
+[5] established well-posedness in real-analytic spaces, whereas Dietert and Gérard-Varet [3] exhibited instability for some smooth initial data, indicating that the analytic class is optimal for well-posedness.
 1.2. Main results
-To simplify the equations (1.10), we introduce the new coordinate $(x,y)$ and new notations $(u,v,P,A)$:
+To simplify the system (1.7), we introduce the new coordinate $(x,y)$ and new unknowns $(u,v,P,A)$:
 (1.10) $x=x_{*},\ y=y_{*}-F(x_{*}),$
 $(u_{*},v_{*},P_{*},A_{*})=(y+u,v+(y+u)\frac{dF}{dx},P,A).$
-Then, the problem (1.10) can be reduced to
+Then the problem (1.7) can be reduced to
 (1.11) $\left\{\begin{aligned} &y\partial_{x}u+v+\partial_{x}|\partial_{x}|A-\partial_{y}^{2}u=-u\partial_{x}u-v\partial_{y}u,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{x}u+\partial_{y}v=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &(u,v)\big|_{y=0}=(0,0),\quad u\big|_{y=f(x)-F(x)}=A(x)+F(x),\quad x\in{\mathbb{R}},\\ &\lim_{x\to-\infty}u=0,\quad 0<y<f(-\infty)-F(-\infty).\end{aligned}\right$.
-For convenience, we take $f(x)=M+F(x)$ for $M=\epsilon^{-\alpha_{0}}$ and $\alpha_{0}\in(0,\frac{3}{4})$. It is useful to introduce the vorticity $\omega=\partial_{y}u$ and
+For convenience, we take $f(x)=M+F(x)$ for $M>0$. We introduce the vorticity $\omega=\partial_{y}u$ and
 (1.12) $u(x,y)=I_{y}[\omega]:=\int_{0}^{y}\omega dy^{\prime}.$
-Differentiating the first equation of (1.13) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows.
+Differentiating the first equation of (1.11) in the variable $y$, we obtain the perturbation system about $(\omega,A)$ as follows
 (1.13) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-Here $F$ is a prescribed wall disturbance function. We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$.
-To establish a rigorous framework, we introduce some functional spaces in this paper. Define the weighted Sobolev spaces for vorticity $\omega$ as
-(1.16) $X_{\bar{\alpha}}=$ $\ \bigg{\{}f\big{|}\ {\big{\|}f\big{\|}}_{X_{\bar{\alpha}}}={\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}<\infty\bigg{\}},$
+Here $F$ is a prescribed function.
+We denote $\Omega_{M}={\mathbb{R}}\times[0,M]$.
+It is easy to see that the Couette flow $(u_{*},v_{*},P_{*},A_{*})$ corresponds to the trivial solution $(u,v,P,A)=(0,0,0,0)$ when $F=0$.
+Now we introduce the functional framework of classical solutions. For the vorticity $\omega$, we define the weighted Sobolev spaces by
+$X_{\bar{\alpha}}=$ $\ \bigg\{f\big|\ {\big\|f\big\|}_{X_{\bar{\alpha}}}={\big\|f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|f\big\|}_{Y_{\bar{\alpha}}}<\infty\bigg\},$
 $Y_{\bar{\alpha}}=$ $\ \bigg\{f\big|\ {\big\|f\big\|}_{Y_{\bar{\alpha}}}={\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big\|}_{L^{2}(\Omega_{M})}$
@@
 $\qquad\qquad\qquad\quad+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}<\infty\bigg\},$
-where $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$.
+where the parameter $\bar{\alpha}$ is nonnegative and $|\partial_{x}|^{s}:=\mathcal{F}^{-1}[|\xi|^{s}\mathcal{F}[\cdot]]$ denotes the Fourier multiplier of order $s$.
 For the displacement function $A$, we define
-(1.17) $X_{\alpha,\infty}=\bigg{\{}f\big{|}\ {\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
-The combined space for $(\omega,A)$ incorporates the upper boundary condition is
-(1.18) ${\mathcal{X}}_{\alpha,M}=\bigg{\{}(\omega,A)\big{|}$ $\ {\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}={\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A\big{\|}}_{X_{\alpha,\infty}}<\infty,$
+$X_{\alpha,\infty}=\bigg\{f\big|\ {\big\||\partial_{x}|^{\frac{5}{6}}f\big\|}_{H^{\frac{4}{3}+\frac{\alpha}{2}}({\mathbb{R}})}+{\big\|f\big\|}_{L^{\infty}({\mathbb{R}})}<\infty\bigg\}.$
+Then the functional space for $(\omega,A)$ incorporating the upper boundary condition is defined by
+${\mathcal{X}}_{\alpha,M}=\bigg\{(\omega,A)\big|$ $\ {\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}={\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|A\big\|}_{X_{\alpha,\infty}}<\infty,$
 $\ \ \omega(\cdot,y)\big|_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg\}.$
 Here $Ai(z)$ is the Airy function satisfying $\frac{d^{2}}{dz^{2}}Ai(z)-zAi(z)=0,\ Ai\big|_{z\uparrow+\infty}=0$.
-We adopt standard mixed-norm notations on $\Omega_{M}$ as following:
-(1.19) $\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big{\|}f\big{\|}}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
-${\big{\|}f\big{\|}}_{L_{x}^{p}L_{y}^{q}}={\big{\|}f\big{\|}}_{L_{x}^{p}({\mathbb{R}};L_{y}^{q}([0,M]))},\quad\mbox{for }\ p,q\geq 1,$
-${\big{\|}f\big{\|}}_{H_{x}^{s}L_{y}^{2}}={\big{\|}f\big{\|}}_{H_{x}^{s}({\mathbb{R}};L_{y}^{2}([0,M]))},\quad{\big{\|}f\big{\|}}_{L_{x}^{2}H_{y}^{s}}={\big{\|}f\big{\|}}_{L_{x}^{2}({\mathbb{R}};H_{y}^{s}([0,M]))},\quad\mbox{for }\ s\in{\mathbb{R}}.$
-For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$, such that $B_{0}\leq CB_{1}$.
-To seek the solution of the system (1.15), we use the following decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ where $(\omega_{0},A_{0})$ satisfies the following linear part of (1.15):
-(1.20) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in{\mathbb{R}}\times[0,M],\\ &\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x),\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-Our first result is to construct a solution of (1.20).
-Theorem 1.1.
-Let $\alpha\in(0,\frac{7}{3}]$, then the linearized system (1.20) has a solutions $(\omega_{0},A_{0})$ belongs to ${\mathcal{X}}_{\alpha,M}$ and satisfies the following estimates:
-(1.21) ${\big{\|}\omega_{0}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A_{0}\big{\|}}_{X_{\alpha,\infty}}\leq$ $\ C{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})},$
-where $C>0$ is independent of $\alpha,M$, and $F$.
-Now, we are the position to state the main results of this paper.
-Theorem 1.2.
-Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$.
-Then, there exist a small enough constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and large enough constant $M_{0}$, such that when $M\geq M_{0}$ and
-${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\delta_{0}M^{-\frac{5+3\varepsilon}{2}},$
-the system (1.15) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ where $(\omega_{0},A_{0})$ defined by (1.20) and $(\bar{\omega},\bar{A})$ satisfying
-(1.22) ${\big{\|}\bar{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}\big{\|}}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2},$
-where $C>0$ is independent of $\alpha,\varepsilon,M$, and $F$. Moreover, $(\omega,A)$ is the unique solution of (1.15) within ${\mathcal{X}}_{\alpha,M}$.
-Remark 1.1.
-Our theoretical results align closely with physical phenomena.
-The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$ under small perturbations $F$, consistent with the principle that linear approximations govern near-equilibrium states in fluid dynamics.
-Remark 1.2.
-Considering the scaling relation $M=O(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$, our analysis shows that the allowable perturbation scale $O(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
-Remark 1.3.
-When $F=0$, our results imply that there is a unique zero solution in ${\mathcal{X}}_{\alpha,M}$ which aligns with the results by Iyer and Maekawa in [9].
-Remark 1.4.
-Here, we take $M\sim\epsilon^{-\alpha_{0}}$ with some small data $\epsilon$ which implies that $M$ will be large enough.
-Thus, taking $M_{0}$ large enough is satisfied by our model.
-On the other hand, for general $M_{0}$, the results may also be got by Remark 3.2.
-1.3. Main ideas
-In this subsection, we state the main ideas of the paper.
-First, we notice that the system (1.15) is a degenerate nonlinear parabolic equation with variable coefficients, where the degeneration occurs on the wall $y=0$.
-To deal with the degeneration, we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ satisfying
+To seek the solution to the system (1.13), we introduce the following crucial decomposition $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$, where $(\omega_{0},A_{0})$ and $(\bar{\omega},\bar{A})$ solve the following system
 (1.14) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}+F,\quad x\in{\mathbb{R}},\end{aligned}\right$.
@@
 (1.15) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(u\partial_{x}\omega+v\partial_{y}\omega)\triangleq{\bar{h}},\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A},\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-It separates the main effects caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
-In the homogeneous problem (1.23), we observe that the integral condition exhibits nonlocal behavior, so it cannot be treated in isolation.
-Instead, a unified framework is needed to couple this integral condition with the boundary data at $y=0$ (i.e., $\partial_{y}\omega_{0}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A_{0}$).
-Their synthesis yields an elliptic multiplier $m(D)^{-1}$ (see (3.45)).
-This multiplier crucially links the displacement $A_{0}$ and vorticity $\omega_{0}$ by
-(1.25) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\lesssim{\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{L^{2}}\lesssim{\big{\|}F\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})},\ \bar{\alpha}\geq 0,$
-providing a key tool for analyzing the well-posedness of the pair $(\omega,A)$.
-In order to solve the non-homogeneous problem (1.24), the nonlinear terms and boundary terms are treated separately. Then, we decompose $\bar{\omega}=\bar{\omega}_{e}+\bar{\omega}_{b}$ satisfying
-(1.26) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e}-\partial_{y}^{2}\bar{\omega}_{e}=-(u\partial_{x}\omega+v\partial_{y}\omega),\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{e}\big{|}_{y=0}=0,\quad x\in{\mathbb{R}},\end{aligned}\right$.
-and
-(1.27) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b}-\partial_{y}^{2}\bar{\omega}_{b}=0,\quad(x,y)\in\Omega_{M}\\ &\partial_{y}\bar{\omega}_{b}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}_{b}]=\bar{A}-I_{M}[\bar{\omega}_{e}],\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-Similar to the homogeneous problem (1.23), we can find a solution $(\bar{\omega}_{b},\bar{A})$ to (1.27) with the following estimates
-(1.28) ${\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{\bar{\alpha}}}\lesssim$ $\ {\big{\|}m(D)^{-1}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}\bar{A}\big{\|}}_{L^{2}}\lesssim{\big{\|}I_{M}[\bar{\omega}_{e}]\big{\|}}_{\dot{H}^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}({\mathbb{R}})}\quad\text{for }\ \bar{\alpha}\geq 0$
-[…]
-Hence, the essential step is to solve (1.26). To this end, we firstly omit the integral condition and consider a linear problem as following
+Our first main result is the solvability of the linear system (1.14).
+Theorem 1.1.
+Let $\alpha\in(0,\frac{7}{3}]$. The linear system (1.14) admits a solution $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$ satisfying the estimate
+(1.16) ${\big\|\omega_{0}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|A_{0}\big\|}_{X_{\alpha,\infty}}\leq$ $\ C{\big\|F\big\|}_{H^{2}({\mathbb{R}})},$
+where $C>0$ is independent of $\alpha,M$.
+The second main result of this paper is the existence and uniqueness of classical solution to the nonlinear system (1.13).
+Theorem 1.2.
+Let $\alpha\in(\frac{2}{3},\frac{7}{3}]$ and $\varepsilon\in(0,\frac{1}{6}]$.
+There exist a small constant $\delta_{0}=\delta_{0}(\alpha,\varepsilon)>0$ and a sufficiently large constant $M_{0}$, such that when $M\geq M_{0}$ and
+${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\delta_{0}M^{-\frac{5+3\varepsilon}{2}},$
+the system (1.13) admits a solution $(\omega,A)=(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}$ with
+${\big\|\bar{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}\big\|}_{X_{\alpha,\infty}}\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
+Here $C>0$ is independent of $\alpha,\varepsilon,M$. Moreover, $(\omega,A)$ is the unique solution to the system (1.13) within ${\mathcal{X}}_{\alpha,M}$ if
+${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq 2\delta_{0}M^{-\frac{5+3\varepsilon}{2}}.$
+Remark 1.1.
+Our theoretical results align closely with the physical phenomena.
+The linear component $(\omega_{0},A_{0})$ captures the dominant behavior of $(\omega,A)$.
+Considering the scaling relation $M={\mathcal{O}}(\text{Re}^{\frac{\alpha_{0}}{8}})$ for $\alpha_{0}\in(0,\frac{3}{4})$ (see Appendix for details), our analysis shows that the allowable roughness scale ${\mathcal{O}}(\text{Re}^{-a_{*}})$ with $a_{*}\in(0,\infty)$ falls within the physically reasonable range $0,\infty)$ (see [26]).
+Remark 1.2.
+If $F=0$, Iyer and Maekawa [9] derived the local rigidity of Couette flow.
+Specifically, they proved that an $L^{2}$ strong solution $(\omega,A)$ of (1.13) is trivial if ${\big\|(\omega,A)\big\|}_{sc}$ is sufficiently small, where ${\big\|\cdot\big\|}_{sc}$ is a scaling invariant norm.
+As a byproduct of Theorem 1.2, we further establish the local uniqueness of Couette flow: the system (1.13) has a locally unique solution in ${\mathcal{X}}_{\alpha,M}$, which is a trivial solution.
+Remark 1.3.
+We adopt $M\gg 1$ in our model for consistency with physical observations. Indeed, for general $M>0$, the results may also be got by Remark 3.2.
+1.3. Main difficulties and Key ideas
+The system (1.13) is a degenerate nonlinear parabolic equation with variable coefficients and its degeneration occurs on the wall $y=0$.
+To solve the system (1.13), we first introduce a crucial decomposition (1.14) and (1.15), which separates the confluence caused by the roughness $F$ from the mathematical difficulties arising from nonlinear terms.
+For the homogeneous problem (1.14), the key ingredient is to solve the following linear problem
 (1.17) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big|_{y=0}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-Upon Fourier transformation, the linearized operator of the first equation in (1.29) is expressed as $i\xi y-\partial_{y}^{2}$, and it shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$.
-Then it is straight to observe that there exists a solution $\hat{f}_{e}$ to the following linearized problem
-$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big{|}_{y=0}=0,\quad\xi\in{\mathbb{R}},\end{aligned}\right$.
-for any given function $h\in L^{2}(\Omega_{M})$, and $\hat{f}$ is the Fourier transformation of $f$ in the variable $x$. Moreover, similar to the proof in [9], the solution satisfies
-(1.30) ${\big{\|}f_{e}\big{\|}}_{Y_{0}}\leq C{\big{\|}h\big{\|}}_{L^{2}}.$
-Here $C>0$ is independent of $M$.
-Thus, in order to solve (1.26), we need to obtain the $L^{2}$-estimate of the nonlinear terms $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$.
-By calculations, we obtain
-(1.31) ${\big{\|}u\partial_{x}\omega+v\partial_{y}\omega\big{\|}}_{L^{2}}\lesssim({\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}+{\big{\|}y^{\frac{1}{2}}\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}+{\big{\|}y\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}^{2}$
-$\qquad\qquad\qquad\qquad\quad+{\big{\|}\bar{\omega}_{b}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data}.$
-Clearly, the norms ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}$ and ${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$ can not be controlled by ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}$.
-Even if we add low-frequency information, the subspace $L^{2}(\Omega_{M})\cap L_{x}^{2}\dot{H}_{y}^{2}\cap\dot{H}_{x}^{\frac{2}{3}}L_{y}^{2}$ of $Y_{0}$ is unable to be embedded into $L^{\infty}$.
-Thus, we need not only a low-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}(\Omega)}$, but also a higher-frequency estimate ${\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}$ for $\alpha>0$ such that
-(1.32) ${\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{L^{2}\cap Y_{\alpha}}^{1-\theta_{1}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{0}}^{\theta_{1}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{X_{\alpha}}^{1-\theta_{1}},\ \text{for some }\theta_{1}\in(0,1),$
-${\big{\|}(y^{\frac{1}{2}}+y)\partial_{y}\bar{\omega}_{e}\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}\lesssim{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{0}}^{\theta_{2}}{\big{\|}\bar{\omega}_{e}\big{\|}}_{Y_{\alpha}}^{1-\theta_{2}},\ \text{for some }\theta_{2}\in(0,1).$
-The main difficulty of this paper is to obtain the low-frequency estimate of $\bar{\omega}_{e}$.
-We observe that the homogeneous solution of the linearized operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$.
-For any nonhomogeneous term $h$, the general solution includes a term of the form $(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz$, where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function.
+Taking the Fourier transformation in $x$, the linear operator of the first equation in (1.17) is expressed as $i\xi y-\partial_{y}^{2}$.
+It shares a structural similarity with the classical Airy operator $\frac{d^{2}}{dz^{2}}-z$.
+Indeed, there exists a solution $\hat{f}_{e}$ to the following linear problem
+$\left\{\begin{aligned} &(i\xi y-\partial_{y}^{2})\hat{f}=\hat{h},\quad(\xi,y)\in\Omega_{M},\\ &\partial_{y}\hat{f}\big|_{y=0}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
+Similar to the proof in [9], the solution satisfies
+${\big\|f_{e}\big\|}_{Y_{0}}\leq C{\big\|h\big\|}_{L^{2}}.$
+However, for the existence of the solution, we need to establish the following $L^{2}$ boundedness
+${\big\|f_{e}\big\|}_{L^{2}}\lesssim{\big\|h\big\|}_{L^{2}},$
+which is the key difference and difficulty compared with the work [9].
+We observe that the homogeneous solution to the linear operator involves a factor $(i\xi)^{-\frac{1}{3}}$, which exhibits a singularity near $\xi=0$.
+For any inhomogeneous term $h$, the general solution includes a term of the form
+$(i\xi)^{-\frac{1}{3}}\int G(y,z;\xi)\hat{h}(\xi,z)\,dz,$
+where the quantity $(i\xi)^{-\frac{1}{3}}G(y,z;\xi)$ represents a specific Green’s function.
 Note that the nonlinear term $u\partial_{x}\omega-v\partial_{y}\omega$ cannot be entirely recast solely as $x$-directional derivatives (i.e., $\partial_{x}(\text{Terms})$).
-This structural constraint hinders low-frequency estimates derived from the nonlinear terms.
+This structural constraint hinders low-order estimates derived from the nonlinear terms.
 To address the singularity near $\xi=0$, a key observation is that the function
@@
 This property is non-trivial (see Remark 3.1), which involves the distribution of zeros of the Bessel function.
-Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form:
-(1.33) $G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
+Based on this property, we find a subtle function $G$ to create $(i\xi)^{\frac{1}{3}}$, which takes the following form
+$G(y,z;\xi)=\frac{\pi}{t((i\xi)^{\frac{1}{3}}M)}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
 $\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y),$
 $\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
-It is straight to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\}$.
-With these observations, we derive that
-(1.34) ${\big{\|}f_{e}\big{\|}}_{L^{2}}\lesssim{\big{\|}h\big{\|}}_{L^{2}}.$
-To control the higher-frequency of $f_{e}$, we introduce the space $Y_{\alpha}$ for $\alpha>0$. It requires higher regularity of $h$ since
-(1.35) ${\big{\|}f_{e}\big{\|}}_{Y_{\alpha}}\lesssim{\big{\|}h\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}},\quad\forall\alpha>0.$
-By product estimates, we finally control the $H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$-norm of nonlinear terms in $\bar{h}$} for $\alpha>\frac{2}{3}$, that is,
-${\big{\|}{\bar{h}}\big{\|}}_{H_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},\quad\forall\alpha>\frac{2}{3}.$
-It follows that
-(1.36) ${\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}\lesssim{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+\text{Norms of given data},\quad\forall\alpha>\frac{2}{3}.$
-Although the above process is sufficient to obtain a solution $(\omega,A)$ in $L^{2}\cap Y_{0}\cap Y_{\alpha}$ for $\alpha>\frac{2}{3}$, it is natural to require that the displacement $A$ is bounded for all $M\geq M_{0}$.
-We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then it satisfies
+It is easy to verify that $\mu_{M}$ eliminates the singularity near $\xi=0$ and $G$ is bounded on $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)\in[-\frac{\pi}{6},\frac{\pi}{6}]\big\}$.
+The last key point is to show that the displacement $A$ is bounded for all $M\geq M_{0}$. We observe that $A=\int_{0}^{M}\omega(x,y)dy$ and then
 ${\big\|A\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}.$
-Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ estimate of $\omega$ via the norm ${\big{\|}\omega\big{\|}}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$.
-To this end, we utilize the interpolation inequality.
-Let $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$ be an interpolation exponent, then we have
+Indeed, we can bound the $L_{x}^{\infty}L_{y}^{1}$ norm of $\omega$ by means of the norm ${\big\|\omega\big\|}_{L^{2}\cap Y_{\bar{\alpha}}}$ for $\bar{\alpha}\geq 0$ via interpolation: there exists some $\theta_{1}=\theta_{1}(\bar{\alpha})\in(0,1)$,
 ${\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq C(\bar{\alpha},M){\big\|\omega\big\|}_{L^{2}}^{\theta_{1}}{\big\|\omega\big\|}_{Y_{0}}^{1-\theta_{1}}+{\big\|\omega\big\|}_{L_{x}^{\infty}({\mathbb{R}};L_{y}^{1}([0,1])}\leq C_{0}(\bar{\alpha},M){\big\|F\big\|}_{H^{2}({\mathbb{R}})}.$
@@
 (1.18) $\inf_{\bar{\alpha}\in[0,\frac{7}{3}]}C_{0}(\bar{\alpha},M)=C_{0}(0,M)=O([\log(M)]^{\frac{1}{2}}),\quad\forall M\geq M_{0}.$
-It implies that the bound of $A$ can not be controlled uniformly in $M$ by the norm ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}$, which contradicts the physical observations that the linear part $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ determines the behavior of $(\omega,A)$.
-To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big{\|}}_{L^{2}}$, so as to derive a refined estimate.
-Specifically, we obtain
+This means that the bound of $A$ cannot be controlled uniformly in $M$ by the norm ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}$—a contradiction to the physical intuition that that the linear component $(\omega_{0},A_{0})$ of the solution $(\omega,A)$ dictates the behavior of $(\omega,A)$.
+To eliminate the dependence of the constant $C_{0}$ on $M$, we introduce a new norm ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}$, which allows us to derive a refined estimate.
+Specifically, we have
 ${\big\|\omega\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq C\big({\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}\omega\big\|}_{L^{2}}+{\big\|\omega\big\|}_{Y_{0}}\big)\leq C{\big\|F\big\|}_{H^{2}({\mathbb{R}})}.$
-Here $C>0$ is independent of $M$. And the estimate of the new norm is derived by using more information of Airy functions (see Proposition 2.3).
-1.4. Organization of the paper
+Here $C>0$ is independent of $M$. The estimate for this new norm is established by leveraging additional properties of Airy functions (see Proposition 2.1).
+1.4. Notations and Organization of the paper
+We adopt standard mixed-norms on $\Omega_{M}$ as follows
+$\langle f,g\rangle_{L^{2}(\Omega_{M})}=\int_{\Omega_{M}}f\bar{g}\ d{\mathbf{x}},\quad{\big\|f\big\|}_{L^{2}}^{2}=\int_{\Omega_{M}}|f|^{2}d{\mathbf{x}},$
+${\big\|f\big\|}_{L_{x}^{p}L_{y}^{q}}={\big\|f\big\|}_{L_{x}^{p}({\mathbb{R}};L_{y}^{q}([0,M]))},\quad\mbox{for }\ p,q\geq 1,$
+${\big\|f\big\|}_{H_{x}^{s}L_{y}^{2}}={\big\|f\big\|}_{H_{x}^{s}({\mathbb{R}};L_{y}^{2}([0,M]))},\quad{\big\|f\big\|}_{L_{x}^{2}H_{y}^{s}}={\big\|f\big\|}_{L_{x}^{2}({\mathbb{R}};H_{y}^{s}([0,M]))},\quad\mbox{for }\ s\in{\mathbb{R}}.$
+For some inequalities, the notation $B_{0}\lesssim B_{1}$ denotes that there exists some $C>0$ independent of $B_{0}$ and $B_{1}$ such that $B_{0}\leq CB_{1}$.
 This paper is organized as follows.
-First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key a priori estimates, including weighted Sobolev norms for vorticity and displacement.
-Subsequently, Section 3 focuses on the linearized problem of the triple-deck system.
-By decomposing the linearized equations into well-posed subproblems (e.g., boundary-value problems for the Airy equation), we construct solutions to the linearized system and establish quantitative bounds on their regularity.
-Thereafter, in Section 4, we apply these linear estimates to prove the main result (Theorem 1.2) via a fixed-point iteration scheme, establishing the existence and uniqueness of solutions to the full nonlinear triple-deck equations.
-Finally, we present an appendix, which provides a rigorous formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer.
-We also present explicit solutions for the upper and main layers, demonstrating their dependence on the viscous sublayer solution.
-This appendix reinforces the self-contained nature of the mathematical framework.
-2. Preliminaries
+First, in Section 2, we characterize the functional spaces used in the analysis of the triple-deck equations and derive key nonlinear estimates, including weighted Sobolev norms for vorticity and displacement.
+Subsequently, Section 3 focuses on the solvability of the linear system of the triple-deck system.
+Thereafter, in Section 4, we prove the main result (Theorem 1.2) via a Picard iteration scheme.
+Finally, we present an appendix, which provides a formal derivation of the complete triple-deck system, including asymptotic matching conditions between the viscous sublayer, main layer, and upper layer.
+2. Functional spaces and nonlinear estimates
 To study the quantity $\omega$, we first consider the space $X_{0}\cap X_{\alpha}$.
-Note that the space $X_{\bar{\alpha}}$ consists of a low-frequency part and a high-frequency part.
-The low-frequency part is characterized by the norms ${\big{\|}f\big{\|}}_{L^{2}(\Omega_{M})}$ and ${\big{\|}|\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big{\|}}_{L^{2}(\Omega_{M})}$, while the high-frequency part is characterized by ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}$.
-We can explore various properties between these two parts.
-First, we present some of them for $f\in X_{0}$.
-Proposition 2.1.
-Assume that $f\in X_{0}$, for any $0\leq\varepsilon\leq\frac{1}{6}$, we have
-(2.1) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}+{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}+{\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}+{\big{\|}f\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big{\|}f\big{\|}}_{X_{0}}.$
-Here $C$ is independent of $\varepsilon$ and $f$.
+Note that the space $X_{\bar{\alpha}}$ consists of a low-order part and a high-order part.
+The low-order part is characterized by the norms ${\big\|f\big\|}_{L^{2}(\Omega_{M})}$ and ${\big\||\partial_{x}|^{\frac{1}{18}}y^{\frac{1}{6}}f\big\|}_{L^{2}(\Omega_{M})}$, while the high-order part is characterized by ${\big\|f\big\|}_{Y_{\bar{\alpha}}}$.
+We now present some properties between these two parts for $f\in X_{0}$.
+Lemma 2.1.
+It holds that for any $0\leq\varepsilon\leq\frac{1}{6}$,
+${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}+{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}+{\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}+{\big\|f\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq C{\big\|f\big\|}_{X_{0}}.$
+Here $C$ is independent of $\varepsilon$.
 Proof.
-By Sobelev embedding theorem and the interpolation theorem, it is straight to obtain
-(2.2) ${\big{\|}f\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}H_{y}^{2}}\leq C{\big{\|}f\big{\|}}_{X_{0}},$
+By Sobelev embedding theorem and the interpolation, we obtain
+${\big\|f\big\|}_{L_{x}^{2}L_{y}^{\infty}}\leq C{\big\|f\big\|}_{L_{x}^{2}H_{y}^{2}}\leq C{\big\|f\big\|}_{X_{0}},$
 ${\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}\leq{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{6\varepsilon}{\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{1-6\varepsilon}\leq{\big\|f\big\|}_{Y_{0}}.$
-Using the inequality ${\big{\|}|\partial_{x}|^{-\frac{1}{3}}f\big{\|}}_{L_{x}^{6}}\leq C{\big{\|}f\big{\|}}_{L_{x}^{2}}$, we derive that
-(2.3) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{6}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}.$
-Having established the above estimates, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through some calculations, we obtain
-(2.4) ${\big{\|}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg{\{}{\big{\|}y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big{\|}y\partial_{x}f\big{\|}}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
+Using the inequality ${\big\||\partial_{x}|^{-\frac{1}{3}}f\big\|}_{L_{x}^{6}}\leq C{\big\|f\big\|}_{L_{x}^{2}}$, we derive
+${\big\|\partial_{y}f\big\|}_{L_{x}^{6}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{1}{3}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}.$
+Having established the estimates above, we now turn to the $L_{y}^{1}L_{x}^{\infty}$-estimate of $f$. Through direct calculations, we obtain
+${\big\|f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq C\min\bigg\{{\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{9}{17}}{\big\|y\partial_{x}f\big\|}_{L_{x}^{2}}^{\frac{8}{17}}\cdot y^{-\frac{19}{34}},$
 $\qquad\qquad\qquad\qquad\quad\ {\big\|y^{\frac{1}{6}}|\partial_{x}|^{\frac{1}{18}}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{3}{11}}{\big\||\partial_{x}|^{\frac{2}{3}}f\big\|}_{L_{x}^{2}}^{\frac{8}{11}}\cdot y^{-\frac{1}{22}}\bigg\},$
@@
 $\qquad\qquad\leq C{\big\|f\big\|}_{X_{0}}.$
-Observe that as the index $\alpha$ increases, higher frequencies in the variable $x$ is accounted for in $X_{\alpha}$, since ${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}f\big{\|}}_{Y_{0}}={\big{\|}f\big{\|}}_{Y_{\alpha}}$.
-These higher frequencies lead to other useful properties.
-Proposition 2.2.
-Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$, then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we obtain
-(2.5) ${\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}+{\big{\|}f\big{\|}}_{L^{\infty}(\Omega_{M})}\leq C\bigg{(}\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\bigg{)}({\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}).$
-Here $C>0$ is independent of $\varepsilon,\alpha$, and $f$.
+Lemma 2.2.
+Assume that $f\in X_{0}\cap X_{\alpha}$ for $\alpha>0$. Then for any $\varepsilon\in(0,\min\{\frac{1}{6},\frac{\alpha}{2}\}]$, we have
+${\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{\infty}}+{\big\|f\big\|}_{L^{\infty}(\Omega_{M})}\leq C\Big(\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}+\frac{1}{\sqrt{\varepsilon}}\Big)\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
+Here $C>0$ is independent of $\varepsilon,\alpha$.
 Proof.
-By interpolation and Proposition 2.1, we have
-(2.6) ${\big{\|}\partial_{y}f\big{\|}}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big{\|}|\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big{\|}|\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big{\|}}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
+By the interpolation and Lemma 2.1, we have
+${\big\|\partial_{y}f\big\|}_{L_{x}^{\infty}({\mathbb{R}})}\leq\frac{C}{\sqrt{\varepsilon}}{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}}{\big\||\partial_{x}|^{\frac{1}{2}+\varepsilon}\partial_{y}f\big\|}_{L_{x}^{2}({\mathbb{R}})}^{\frac{1}{2}},$
 ${\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{\infty}}\leq\frac{C}{\sqrt{\varepsilon}}{\big\||\partial_{x}|^{\frac{1}{2}+\varepsilon}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}{\big\||\partial_{x}|^{\frac{1}{2}-\varepsilon}y^{\frac{1}{2}-3\varepsilon}\partial_{y}f\big\|}_{L_{y}^{2}L_{x}^{2}}^{\frac{1}{2}}$
-$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\bigg{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\bigg{)}{\big{\|}f\big{\|}}_{Y_{0}}^{\frac{1}{2}}$
-$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}({\big{\|}f\big{\|}}_{Y_{0}}+{\big{\|}f\big{\|}}_{Y_{\alpha}}).$
-By classical extension, we denote $\tilde{f}$ as the extension of $f$ to ${\mathbb{R}}^{2}$.
-It is evident to verify that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
-(2.7) ${\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
+$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\Big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}+{\big\||\partial_{x}|^{\frac{1+\alpha}{2}}y^{\frac{1}{2}}\partial_{y}f\big\|}_{L^{2}(\Omega_{M})}^{\frac{1}{2}}\Big){\big\|f\big\|}_{Y_{0}}^{\frac{1}{2}}$
+$\qquad\qquad\qquad\quad\ \leq\frac{C}{\sqrt{\varepsilon}}\big({\big\|f\big\|}_{Y_{0}}+{\big\|f\big\|}_{Y_{\alpha}}\big).$
+We denote by $\tilde{f}$ the extension of $f$ to ${\mathbb{R}}^{2}$ such that $\tilde{f}\in H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))\cap H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))$ for $\alpha>0$ and
+${\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}\leq C{\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))},$
 ${\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\leq C{\big\|f\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}(0,M))}.$
-By calculations, we have
-(2.8) ${\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\hat{\tilde{f}}\big{\|}}_{L_{\xi,\eta}^{1}}$
+Direct calculations show that
+${\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\hat{\tilde{f}}\big\|}_{L_{\xi,\eta}^{1}}$
 $=$ $\ \int_{{\mathbb{R}}^{2}}\bigg\{\big|\hat{\tilde{f}}(\xi,\eta)(1+|\eta|)^{2}\big|^{\theta}\big|\hat{\tilde{f}}(\xi,\eta)(1+|\xi|)^{\frac{2}{3}+\frac{\alpha}{2}}\big|^{1-\theta}$
@@
 $\leq$ $\ C(\alpha,\theta){\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}^{\theta}{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}^{1-\theta}$
-$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}.$
-Here $C(\alpha,\theta)={\big{\|}(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big{\|}}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$ in the last inequality.
-It implies that
-(2.9) ${\big{\|}f\big{\|}}_{L_{x,y}^{\infty}}\leq$ $\ {\big{\|}\tilde{f}\big{\|}}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}\tilde{f}\big{\|}}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big{\|}\tilde{f}\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\bigg{)}$
-$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big{\|}f\big{\|}}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big{(}{\mathbb{R}};L_{y}^{2}((0,M))\big{)}}\bigg{)}$
-$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\bigg{(}{\big{\|}f\big{\|}}_{X_{0}}+{\big{\|}f\big{\|}}_{X_{\alpha}}\bigg{)}.$
-Next, we present some a priori estimates for the nonlinear terms. Suppose $(u,v)$ satisfies
-$\partial_{x}u+\partial_{y}v=0,$
-with the boundary data
+$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big).$
+Here $C(\alpha,\theta)={\big\|(1+|\eta|)^{-2\theta}(1+|\xi|)^{-(\frac{2}{3}+\frac{\alpha}{2})(1-\theta)}\big\|}_{L_{\xi,\eta}^{2}}$ and $\theta=\frac{1}{2}(\frac{1}{4}+\frac{1+3\alpha}{4+3\alpha})$.
+This shows that
+${\big\|f\big\|}_{L_{x,y}^{\infty}}\leq$ $\ {\big\|\tilde{f}\big\|}_{L_{x,y}^{\infty}}\leq C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|\tilde{f}\big\|}_{H_{y}^{2}({\mathbb{R}};L_{x}^{2}({\mathbb{R}}))}+{\big\|\tilde{f}\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}({\mathbb{R}};L_{y}^{2}({\mathbb{R}}))}\big)$
+$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{H_{y}^{2}((0,M);L_{x}^{2}({\mathbb{R}}))}+{\big\|f\big\|}_{H_{x}^{\frac{2}{3}+\frac{\alpha}{2}}\big({\mathbb{R}};L_{y}^{2}((0,M))\big)}\big)$
+$\leq$ $\ C\frac{(1+\alpha)^{\frac{1}{2}}}{\alpha}\big({\big\|f\big\|}_{X_{0}}+{\big\|f\big\|}_{X_{\alpha}}\big).$
+Next we are devoted to the estimates of nonlinear terms. Let the velocity $(u,v)$ satisfy the divergence-free condition
+$\partial_{x}u+\partial_{y}v=0$
+subject to the boundary conditions
 $(u,v)\big|_{y=0}=0,\quad u\big|_{y=M}=A_{M}.$
-Set $h=-(u\partial_{x}\omega+v\partial_{y}\omega)$ for $\omega=\partial_{y}u$, then we have
-(2.10) $h=-u\partial_{x}\omega-\bigg{(}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz-y\partial_{x}A_{M}\bigg{)}\partial_{y}\omega.$
-Define
+where $A_{M}$ denotes a fixed velocity at the upper boundary $y=M$. Define the function
+$h=-(u\partial_{x}\omega+v\partial_{y}\omega),$
+for $\omega=\partial_{y}u$. Using the divergence-free condition and boundary conditions, we can rewrite $h$ as
+$h=-u\partial_{x}\omega-(V^{0}-y\partial_{x}A_{M})\partial_{y}\omega.$
+Here
 (2.1) $V^{0}=\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz.$
-Next, we give the estimates for the nonlinear term $h$.
 Proposition 2.1.
-Fixing $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$, suppose $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$, then we have $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$.
-Moreover, there holds
-(2.12) ${\big{\|}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},\quad\forall\ \varepsilon\in(0,\frac{1}{6}]$
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-Here $C=C(\alpha,\varepsilon)>0$ is independent of $M,\ \omega$, and $A_{M}$.
+Let $\alpha>\frac{2}{3},\ \varepsilon\in(0,\frac{1}{6}]$.
+Assume that $\omega\in X_{0}\cap X_{\alpha}$ and $\partial_{x}A_{M}\in L^{\infty}\cap H^{\frac{\alpha}{2}}$.
+Then $h\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ with
+(2.2) ${\big\|h\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
+${\big\||\partial_{x}|^{\frac{\alpha}{2}}h\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
+Here $C=C(\alpha,\varepsilon)>0$ is independent of $M$ and $A_{M}$.
 Proof.
-In view of Proposition 2.1 and Proposition 2.2, it is straight to obtain
+By virtue of Lemma 2.1 and Lemma 2.2, we obtain
 ${\big\|h\big\|}_{L^{2}(\Omega_{M})}\leq$ $\ C{\big\|\frac{u}{y}\big\|}_{L^{\infty}(\Omega_{M})}{\big\|y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}+CM^{\frac{3\varepsilon}{2}}{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
 $\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
-$\leq$ $\ C({\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}){\big{\|}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}$
+$\leq$ $\ C\big({\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}+M^{\frac{3\varepsilon}{2}}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}\big){\big\|y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}$
 $\ \ +CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
-$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-which implies the first desired result.
-For the high-order estimates, by product estimates, we have
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
-$\qquad\leq C{\big{\|}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})}$
-$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big{\|}y|\partial_{x}|^{\frac{1}{3}}\omega\big{\|}}_{L^{\infty}(\Omega_{M})},$
+$\leq$ $\ C(1+M^{\frac{1+3\varepsilon}{2}}){\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+CM^{\frac{1+3\varepsilon}{2}}{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
+which gives the first estimate of (2.2).
+For the higher-order estimates, by product estimates, we have
+(2.3) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}$
+$\leq C{\big\|\omega\big\|}_{L^{\infty}(\Omega_{M})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})}$
+$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}{\big\|y|\partial_{x}|^{\frac{1}{3}}\omega\big\|}_{L^{\infty}(\Omega_{M})},$
 and
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big{\|}}_{L^{2}(\Omega_{M})}$
-$\qquad\leq C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{2}L_{y}^{\infty}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}$
-$\qquad\qquad+C{\big{\|}y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big{\|}}_{L_{x}^{4}L_{y}^{\infty}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
-$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big{\|}}_{L^{2}(\Omega_{M})}{\big{\|}y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big{\|}}_{L_{y}^{2}L_{x}^{\infty}}$
-$\qquad\qquad+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}$
-$\qquad\leq CM^{\frac{3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+C{\big{\|}\tilde{z}\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}.$
-All we left is to give the estimates of the right hand of above inequality. By Sobolev embedding Theorem, we have
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+C{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
+(2.4) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}$
+$\leq C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{2}L_{y}^{\infty}}{\big\|y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}$
+$\qquad+C{\big\|y^{-\frac{1}{2}}\int_{0}^{y}\int_{z}^{M}\partial_{x}\omega(\cdot,\tilde{z})d\tilde{z}dz\big\|}_{L_{x}^{4}L_{y}^{\infty}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
+$\leq CM^{\frac{3\varepsilon}{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}\tilde{z}\partial_{x}\omega\big\|}_{L^{2}(\Omega_{M})}{\big\|y^{\frac{1-3\varepsilon}{2}}\partial_{y}\omega\big\|}_{L_{y}^{2}L_{x}^{\infty}}$
+$\qquad+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$
+$\leq CM^{\frac{3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}+C{\big\|\tilde{z}\partial_{x}\omega\big\|}_{L_{x}^{4}L_{\tilde{z}}^{2}}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}.$
+It remains to estimate the right-hand side of the above inequality. By the Sobolev embedding, we have
+(2.5) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{H_{x}^{\frac{2}{3}}L_{y}^{2}}\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+C{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{2}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
 $\qquad\leq C{\big\|\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\alpha}{3}}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\leq C{\big\|\omega\big\|}_{X_{\alpha}},$
@@
 and
-${\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{4}L_{y}^{2}}\leq C{\big{\|}|\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
-$\qquad\leq C\big{(}{\big{\|}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
-$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}},$
+(2.6) ${\big\|y\partial_{x}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$ $\leq C{\big\||\partial_{x}|^{\frac{1}{4}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
+$\leq C\big({\big\|y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{x}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)\quad\mbox{for}\ \ \alpha\geq\frac{1}{2}$
+$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}},$
 ${\big\||\partial_{x}|^{\frac{\alpha}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{4}L_{y}^{2}}$ $\leq{\big\||\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{4}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
-$\qquad\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}\big{)}$
-$\qquad\leq C{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-Combining all above estimates, for $\alpha>\frac{2}{3}$, we obtain
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big{\|}}_{L^{2}(\Omega_{M})}\leq$ $\ CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}^{2},$
+$\leq C\big({\big\||\partial_{x}|^{\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}+\frac{1}{2}}y^{\frac{1}{2}}\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}\big)$
+$\leq C{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}.$
+Combining all estimates above, we deduce that for $\alpha>\frac{2}{3}$,
+(2.7) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(u\partial_{x}\omega)\big\|}_{L^{2}(\Omega_{M})}\leq\ CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2},$
 ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(V^{0}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}\leq\ CM^{\frac{3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}^{2}.$
@@
 (2.8) ${\big\||\partial_{x}|^{\frac{\alpha}{2}}(y\partial_{x}A_{M}\partial_{y}\omega)\big\|}_{L^{2}(\Omega_{M})}$
-$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}{\big{\|}y\partial_{y}\omega\big{\|}}_{L_{x}^{\infty}L_{y}^{2}}+C{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})}{\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big{\|}}_{L_{x}^{2}L_{y}^{2}}$
-$\qquad\leq CM^{\frac{1+3\varepsilon}{2}}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}}$
-$\qquad\leq CM{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
+$\leq C{\big\||\partial_{x}|^{\frac{\alpha}{2}}\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}{\big\|y\partial_{y}\omega\big\|}_{L_{x}^{\infty}L_{y}^{2}}+C{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})}{\big\||\partial_{x}|^{\frac{\alpha}{2}}y\partial_{y}\omega\big\|}_{L_{x}^{2}L_{y}^{2}}$
+$\leq CM^{\frac{1+3\varepsilon}{2}}{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{\dot{H}^{\frac{\alpha}{2}}}+CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}}$
+$\leq CM{\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}.$
+Substituting (2.5)-(2.8) into (2.3)-(2.4), we conclude the last estimate of (2.2). ∎
 Remark 2.1.
 Indeed, if $h$ is replaced by $h_{1}=-(u\partial_{x}\omega_{1}+v\partial_{y}\omega_{1})$ for $\omega_{1}\in X_{0}\cap X_{\alpha}$, then by repeating the proof of Proposition 2.1, we have
-(2.14) ${\big{\|}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}},\ \ \forall\varepsilon\in(0,\frac{1}{6}]$
-${\big{\|}|\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big{\|}}_{L^{2}(\Omega_{M})}\leq CM\bigg{(}{\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\partial_{x}A_{M}\big{\|}}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\bigg{)}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-Next we give priori estimate of $I_{M}[f]$ for $f\in X_{\bar{\alpha}}$.
-Proposition 2.4.
-Suppose $\bar{\alpha}\geq 0$, for any $f\in X_{\bar{\alpha}}$, we have
+${\big\|h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{2}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}},\ \ \varepsilon\in(0,\frac{1}{6}],$
+${\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{1}\big\|}_{L^{2}(\Omega_{M})}\leq CM\big({\big\|\omega\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\partial_{x}A_{M}\big\|}_{L^{\infty}({\mathbb{R}})\cap\dot{H}^{\frac{\alpha}{2}}({\mathbb{R}})}\big){\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}.$
+The following lemma is devoted to the integral boundary condition $I_{M}[\omega]=A(x)+F(x)$.
+Lemma 2.3.
+It holds that for any $\bar{\alpha}\geq 0$,
 (2.9) ${\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big\|}_{L^{2}({\mathbb{R}})}\leq$ $\ C{\big\|f\big\|}_{Y_{\bar{\alpha}}}.$
-Here $C$ is independent of $\bar{\alpha}$ and $f$.
+Here $C$ is independent of $\bar{\alpha}$.
 Proof.
 Observe that
-$\big{|}\int_{0}^{M}\hat{f}(\xi,y)dy\big{|}=$ $\ \big{|}\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\big{|}$
+$\left|\int_{0}^{M}\hat{f}(\xi,y)dy\right|=$ $\left|\int_{0}^{|\xi|^{-\frac{1}{3}}}\hat{f}(\xi,y)dy+\int_{|\xi|^{-\frac{1}{3}}}^{M}y^{-1}y\hat{f}(\xi,y)dy\right|$
 $\leq$ $\ C|\xi|^{-\frac{1}{6}}{\big\|\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}}+C|\xi|^{\frac{1}{6}}{\big\|y\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}},$
-which implies that
-(2.16) $\big{|}\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{|}\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big{\|}\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big{\|}y\hat{f}(\xi,\cdot)\big{\|}}_{L_{y}^{2}}).$
-Thus,
-(2.17) ${\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big{\|}}_{L^{2}({\mathbb{R}})}\leq C\big{(}{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L^{2}}+{\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big{\|}}_{L^{2}}\big{)}\leq C{\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}},$
-which implies the desired results. ∎
-3. Proof of Theorem 1.1
-In this section, we give the proof of the Theorem 1.1. Before that, we study a nonhomogeneous linearized problem without the integral condition as follows.
+which implies
+$\big|\xi^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big|\leq$ $\ C|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\cdot{\big\|\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}({\mathbb{R}})}+C|\xi|^{1+\frac{\bar{\alpha}}{2}}{\big\|y\hat{f}(\xi,\cdot)\big\|}_{L_{y}^{2}}).$
+Thus, we obtain
+${\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}I_{M}[f]\big\|}_{L^{2}({\mathbb{R}})}\leq C\big({\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big\|}_{L^{2}}+{\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}f\big\|}_{L^{2}}\big)\leq C{\big\|f\big\|}_{Y_{\bar{\alpha}}}.$
+It completes the proof. ∎
+3. The solvability of the linear system
+Recall that, to solve the nonlinear system (1.13), we decompose $\omega=\omega_{0}+\bar{\omega}$ and $A=A_{0}+\bar{A}$ with
+(3.1) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0},\quad I_{M}[\omega_{0}]=A_{0}+F,\quad x\in{\mathbb{R}},\end{aligned}\right$.
+and
+(3.2) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(u\partial_{x}\omega+v\partial_{y}\omega)\triangleq{\bar{h}},\quad(x,y)\in\Omega_{M},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A},\quad x\in{\mathbb{R}}.\end{aligned}\right$.
+To address the system (3.2), we first solve an inhomogeneous linear problem without the integral condition
 (3.3) $\left\{\begin{aligned} &y\partial_{x}f_{e}-\partial_{y}^{2}f_{e}=h_{e},\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f_{e}\big|_{y=0}=0,\quad f_{e}\big|_{y=M}=0,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-Lemma 3.1.
-Fix $\alpha>0$. We assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$. Then, there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.1). Moreover, there holds
+Proposition 3.1.
+Assume that $h_{e}\in H_{x}^{\frac{\alpha}{2}}L_{y}^{2}$ for some $\alpha>0$. Then there exists a unique solution $f_{e}\in X_{0}\cap X_{\alpha}$ to the problem (3.3). Moreover, there holds that
 ${\big\|f_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq C\big[(1+M^{2}){\big\|h_{e}\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big\|}_{L^{2}(\Omega_{M})}\big],$
-where $C$ is a constant independent of $\alpha,M$, and $h_{e}$.
+where $C$ is a constant independent of $\alpha,M$.
 Proof.
 The proof proceeds in three steps.
-Step 1. We give the explicit form of the solution. Taking the Fourier transform in variable $x$ into (3.1), we obtain
+Step 1. The explicit formula of the solution.
+Taking the Fourier transform in variable $x$ into (3.3), we obtain
 (3.4) $\left\{\begin{aligned} &i\xi y\hat{f}_{e}-\partial_{y}^{2}\hat{f}_{e}=\hat{h}_{e},\quad(\xi,y)\in{\mathbb{R}}\times(0,M),\\ &\partial_{y}\hat{f}_{e}\big|_{y=0}=0,\ \hat{f}_{e}\big|_{y=M}=0,\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
-Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.2), which has an explicit formula as following:
+Using the Green-function method, we find the unique solution $\hat{f}_{e}$ of (3.4), which has an explicit formula
 (3.5) $\hat{f}_{e}(\xi,y)=$ $\ (i\xi)^{-\frac{1}{3}}\int_{0}^{M}G(y,z;\xi)\hat{h}_{e}(\xi,z)dz,$
-with
-(3.4) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M.\end{aligned}\right$.
+with the definition of $G$ given by
+(3.6) $G(y,z;\xi)=\pi s_{M}\left\{\begin{aligned} \mu_{M}(\xi,y)\mu_{0}(\xi,z),&\quad 0\leq z<y,\\ \mu_{M}(\xi,z)\mu_{0}(\xi,y),&\quad y<z\leq M,\end{aligned}\right$.
 $s_{M}=\bigg(\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg)^{-1},$
@@
 $\mu_{0}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}Bi((i\xi)^{\frac{1}{3}}y).$
-Step 2. We derive the low-frequency estimates of $f_{e}$. Since $\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=-\frac{1}{\sqrt{3}}$, we have
-(3.5) $t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
+Step 2. The low-order estimates of $f_{e}$.
+To this end, we first claim that $s_{M}$ is bounded. Since the ratio $\frac{Ai^{\prime}(z)}{Bi^{\prime}(z)}$ of the Airy functions at $z=0$ is $\frac{1}{\sqrt{3}}$, we then define
+$t(z):=\frac{Ai(z)}{Bi(z)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}=\frac{Ai(z)}{Bi(z)}+\frac{1}{\sqrt{3}}=\frac{\sqrt{3}Ai(z)+Bi(z)}{\sqrt{3}Bi(z)}.$
 It is well-known that $Ai(z)$ and $Bi(z)$ are holomorphic on $\mathbb{C}$} and do not vanish for $\text{Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$ [28, Page 15].
@@
 $C_{\theta}=\big\{z\in{\mathbb{C}}\big|\ \text{Arg}(z)\in(-\theta,\theta)\big\}.$
-Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$. Moreover, for any $\delta\in(0,\frac{\pi}{3})$, there holds
+Then, $t(z)$ is a holomorphic and non-vanishing function on $C_{\frac{\pi}{3}}$.
+Moreover, for any $\delta\in(0,\frac{\pi}{3})$, the Airy functions $Ai,Bi$ satisfy the following asymptotic expansions(see [20, Chapter 9]):
 (3.7) $Ai(z)\sim\frac{1}{3^{\frac{2}{3}}\Gamma(\frac{1}{3})}-\frac{1}{3^{\frac{1}{3}}\Gamma(\frac{2}{3})}z,\quad Bi(z)\sim\frac{1}{3^{\frac{1}{6}}\Gamma(\frac{1}{3})}+\frac{3^{\frac{1}{6}}}{\Gamma(\frac{2}{3})}z,\qquad\mbox{as}\ \ z\to 0,$
 $Ai(z)\sim\frac{e^{-\frac{2}{3}z^{\frac{3}{2}}}}{2\sqrt{\pi}z^{\frac{1}{4}}},\quad Bi(z)\sim\frac{e^{\frac{2}{3}z^{\frac{3}{2}}}}{\sqrt{\pi}z^{\frac{1}{4}}},\qquad\mbox{as}\ \ |z|\to\infty\ \ \big(z\in C_{\frac{\pi}{3}-\delta}\big).$
-By (3.6), we have
+By the asymptotic behavior (3.7), we have
 (3.8) $\lim_{|z|\to\infty,\atop|\text{Arg}(z)|\in[0,\frac{\pi}{3}-\delta)}t(z)=\frac{1}{\sqrt{3}}.$
@@
 $|t(z)|\geq c_{0}^{-1},\quad\text{for }z\in\overline{C_{\frac{\pi}{6}}}.$
-Since $s_{M}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we have
+Thanks to $s_{M}=\bigg(\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}-\frac{Ai^{\prime}(0)}{Bi^{\prime}(0)}\bigg)^{-1}=\frac{1}{t((i\xi)^{\frac{1}{3}}M)}$, we then obtain
 (3.9) $|s_{M}|\leq c_{0},\qquad\forall\ \xi\in{\mathbb{R}},\ M>0.$
-Then for $|\xi|\leq M^{-3}$, we obtain
-(3.9) $|\mu_{M}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}$
-$\qquad\qquad\ \leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M],$
-$|\mu_{0}(\xi,y)|=\bigg{|}Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\bigg{|}\leq C,\quad\forall\ y\in[0,M].$
+To derive a low-order estimate of $f_{e}$, we decompose the frequency space into two regimes.
+For the low-frequency regime where $|\xi|\leq M^{-3}$, we observe that $|(i\xi)^{\frac{1}{3}}y|\leq 1$ for all $y\in[0,M]$.
+By the continuity of $Ai$ and $Bi$ on $\mathbb{C}$}, we have
+(3.10) $|\mu_{0}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)+\frac{1}{\sqrt{3}}Bi((i\xi)^{\frac{1}{3}}y)\big|\leq C,\quad\forall\ y\in[0,M].$
+Next, we compute
+(3.11) $\mu_{M}(\xi,y)=Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)$
+$\qquad\quad\ \ =Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}{Bi\big((i\xi)^{\frac{1}{3}}(M-y)+(i\xi)^{\frac{1}{3}}y\big)}Bi((i\xi)^{\frac{1}{3}}y).$
+Since $Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Bi((i\xi)^{\frac{1}{3}}y)}Bi((i\xi)^{\frac{1}{3}}y)=0$ and $Bi(z)\not=0$ for ${\rm Arg}(z)\in(-\frac{\pi}{3},\frac{\pi}{3})$, an application of the Mean Value Theorem to (3.11) yields
+(3.12) $|\mu_{M}(\xi,y)|\leq C|\xi|^{\frac{1}{3}}(M-y),\quad\forall\ y\in[0,M].$
 It follows that
-(3.10) $|\hat{f}_{e}(\xi,y)|\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
+(3.13) $|\hat{f}_{e}(\xi,y)|=$ $\ \bigg|\pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg\{\mu_{M}(\xi,y)\cdot\int_{0}^{y}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)dz$
+$\qquad\qquad\qquad\quad+\mu_{0}(\xi,y)\cdot\int_{y}^{M}\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)dz\bigg\}\bigg|$
+$\leq$ $\ C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
 $\qquad\qquad\qquad\quad+\big|\mu_{0}(\xi,y)\big|\cdot\int_{y}^{M}\big|\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz\bigg\}$
-$\leq$ $\ C(M-y)y^{\frac{1}{2}}{\big{\|}\hat{h}_{e}(\xi,\cdot)\big{\|}}_{L_{z}^{2}([0,y])}$
+[…]
 $\quad+C|\xi|^{-\frac{1}{3}}{\big\||\xi|^{\frac{1}{3}}(M-z)\big\|}_{L_{z}^{2}([y,M])}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([y,M])}$
 $\leq$ $\ C(M-y)M^{\frac{1}{2}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
-Hence, we have
+Hence, we obtain a point-wise estimate of $\hat{f}_{e}$ for the low frequency as follows
 (3.14) ${\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big\|}_{L_{y}^{2}(0,M)}\leq$ $\ CM^{2}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)},$
 for $\alpha^{\prime}\geq 0$.
-And for $|\xi|\geq M^{-3}$, we have
+For the high frequency $|\xi|\geq M^{-3}$, by the asymptotic behavior (3.7) of Airy functions, we have
 (3.15) $|\mu_{M}(\xi,y)|=\big|Ai((i\xi)^{\frac{1}{3}}y)-\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Bi((i\xi)^{\frac{1}{3}}M)}Bi((i\xi)^{\frac{1}{3}}y)\big|$
@@
 $\qquad\qquad\leq C[1+|\xi|^{\frac{1}{3}}y]^{-\frac{1}{4}}e^{\frac{2}{3}{\rm Re\ }[(i\xi)^{\frac{1}{3}}y]^{\frac{3}{2}}},$
-and
+and then,
 (3.16) ${\big\|\mu_{0}(\xi,z)\big\|}_{L_{z}^{2}(0,y)}\leq C|\xi|^{-\frac{1}{6}}\bigg({\big\|Ai(z)\big\|}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}+{\big\|Bi(z)\big\|}_{L_{z}^{2}(0,(i\xi)^{\frac{1}{3}}y)}\bigg)$
@@
 Combining (3.5), (3.6), (3.9), (3.15) with (3.16), we obtain
-(3.14) $|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg{\{}\big{|}\mu_{M}(\xi,y)\big{|}\cdot\int_{0}^{y}\big{|}\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big{|}dz$
+$|\hat{f}_{e}(\xi,y)|\leq C|s_{M}|\cdot|\xi|^{-\frac{1}{3}}\bigg\{\big|\mu_{M}(\xi,y)\big|\cdot\int_{0}^{y}\big|\mu_{0}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz$
 $\qquad\qquad\qquad\qquad\qquad\quad+\big|\mu_{0}(\xi,y)\big|\cdot\int_{y}^{M}\big|\mu_{M}(\xi,z)\hat{h}_{e}(\xi,z)\big|dz\bigg\}$
@@
 $\qquad\leq C|\xi|^{-\frac{1}{2}}[1+|\xi|^{\frac{1}{3}}y]^{-\frac{3}{4}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}([0,M])},\quad\mbox{for}\ \ y\in[0,M].$
-Hence, we have
+Hence, we obtain a pointwise estimate of $\hat{f}_{e}$ for the high frequency as follows
 (3.17) ${\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{f}_{e}(\xi,\cdot)\big\|}_{L_{y}^{2}(0,M)}\leq C(\alpha^{\prime})|\xi|^{-\frac{1}{2}}{\big\|\hat{h}_{e}(\xi,\cdot)\big\|}_{L_{z}^{2}(0,M)}\cdot|\xi|^{-\frac{1}{6}}$
@@
 Here $C$ is independent of $M$ and $f_{e}$.
-Step 3. We derive the high-frequency estimates of $f_{e}$. For any $\bar{\alpha}\geq 0$, we set
+Step 3. The high-order estimates of $f_{e}$.
+For $\bar{\alpha}\geq 0$, we set
 (3.19) $B_{0}^{2}=|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big\|\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}.$
-Multiplying the equation (3.2) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we have
-(3.18) $i|\xi|^{2+\bar{\alpha}}{\big{\|}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
-Here we use a fact that $\hat{f}_{e}\big{|}_{y=M}=0$. Set
+Multiplying the equation (3.4) by $\xi^{1+\bar{\alpha}}y\hat{f}_{e}$, we obtain
+$i|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}=\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}.$
+Here we used the fact that $\hat{f}_{e}\big|_{y=M}=0$. Set
 (3.20) $B_{1}^{2}=$ $\ |\xi|^{1+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2},$
@@
 (3.21) $|\xi|^{2+\bar{\alpha}}{\big\|y\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+B_{1}^{2}$ $\leq\big|-\xi^{1+\bar{\alpha}}\langle\partial_{y}\hat{f}_{e},\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\hat{h}_{e},y\xi^{1+\bar{\alpha}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big|$
-$\qquad\leq\big{|}-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big{|}$
-$\qquad\leq C\bigg{(}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}\bigg{)}B_{0}$
-$\qquad\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big{\|}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}).$
+$\leq\big|-\langle\xi^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e},\xi^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}+\langle\xi^{\frac{\bar{\alpha}}{2}}\hat{h}_{e},y\xi^{1+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\rangle_{L_{y}^{2}}\big|$
+$\leq C\bigg(|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}\bigg)B_{0}$
+$\leq CB_{0}(B_{2}+|\xi|^{\frac{\bar{\alpha}}{2}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}).$
 Here $C$ is independent of $\bar{\alpha}$. Multiplying the equation (3.4) by $|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}$, we obtain
 (3.22) $i\xi|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}=\langle\hat{h}_{e},|\xi|^{\frac{2}{3}+\bar{\alpha}}\hat{f}_{e}\rangle,\quad\mbox{by}\ \ \hat{f}_{e}(\cdot,M)=0.$
-Then we get
-(3.22) $|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big{\|}y^{\frac{1}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big{\|}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}^{2}\leq{\big{\|}|\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}$
-$\qquad\leq CB_{0}{\big{\|}|\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big{\|}}_{L_{y}^{2}}.$
-Thus, by (3.19), we obtain
+Then we have
+$|\xi|^{\frac{5}{3}+\bar{\alpha}}{\big\|y^{\frac{1}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}+|\xi|^{\frac{2}{3}+\bar{\alpha}}{\big\|\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}$ $\leq{\big\||\xi|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}$
+$\leq CB_{0}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}.$
+Thus, we get by (3.20) that
 (3.23) $B_{2}^{2}\leq$ $\ CB_{0}{\big\||\xi|^{\frac{\bar{\alpha}}{2}}\hat{h}_{e}\big\|}_{L_{y}^{2}}.$
-On the other hand, it is straight to obtain
+On the other hand, we have
 (3.24) $|\xi|^{\frac{4}{3}+\bar{\alpha}}{\big\|\hat{f}_{e}\big\|}_{L_{y}^{2}}^{2}$ $=|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}\partial_{y}(y)|\hat{f}_{e}|^{2}dy$
-$\qquad=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big{|}_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
-$\qquad\leq 2{\big{\|}|\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big{\|}}_{L_{y}^{2}}{\big{\|}|\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big{\|}}_{L_{y}^{2}}$
-$\qquad\leq CB_{0}B_{2}.$
-By (3.17), (3.19), (3.20), (3.23), and (3.24), we have
+$=|\xi|^{\frac{4}{3}+\bar{\alpha}}M\hat{f}_{e}^{2}\big|_{y=M}-|\xi|^{\frac{4}{3}+\bar{\alpha}}\int_{0}^{M}y(\bar{\hat{f}}_{e}\partial_{y}\hat{f}_{e}+\hat{f}_{e}\partial_{y}\bar{\hat{f}}_{e})dy$
+$\leq 2{\big\||\xi|^{1+\frac{\bar{\alpha}}{2}}y\hat{f}_{e}\big\|}_{L_{y}^{2}}{\big\||\xi|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\hat{f}_{e}\big\|}_{L_{y}^{2}}$
+$\leq CB_{0}B_{2}.$
+By (3.19), (3.20), (3.21), (3.23) and (3.24), we have
 (3.25) $B_{0}^{2}+B_{1}^{2}\leq C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}.$
-Then by (3.23) and (3.25), there holds
+Then by (3.23) and (3.25), we get
 (3.26) $B_{2}^{2}\leq$ $\ CB_{0}^{2}+C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}\leq C|\xi|^{\bar{\alpha}}{\big\|\hat{h}_{e}\big\|}_{L_{y}^{2}}^{2}.$
-Applying Plancherel Theorem, we obtain
+Applying Plancherel’s theorem, we obtain
 (3.27) ${\big\|f_{e}\big\|}_{Y_{\bar{\alpha}}}^{2}\leq$ $\ C\bigg({\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}y\partial_{x}f_{e}\big\|}_{L^{2}}^{2}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f_{e}\big\|}_{L^{2}}^{2}+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f_{e}\big\|}_{L^{2}}^{2}$
@@
 $=$ $\ C{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}h_{e}\big\|}_{L^{2}}^{2}.$
-By (3.16) and (3.27), we get
+Finally, it follows from (3.18) and (3.27) that
 (3.28) ${\big\|f_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq C\big[(1+M^{2}){\big\|h_{e}\big\|}_{L^{2}}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}h_{e}\big\|}_{L^{2}}\big].$
-Here $C$ is independent of $\alpha,M$, and $f_{e}$.
+Here $C$ is independent of $\alpha,M$, and $f_{e}$. ∎
 Remark 3.1.
-The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows.
-(3.29) $Ai(z)=\frac{\sqrt{z}}{3}\big{(}I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big{)},\quad z\in{\mathbb{C}},$
+The Airy functions $Ai(z),Bi(z)$ admit decompositions [20, Chapter 9] as follows
+$Ai(z)=\frac{\sqrt{z}}{3}\big(I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})-I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big),\quad z\in{\mathbb{C}},$
 $Bi(z)=\sqrt{\frac{z}{3}}\big(I_{-\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})+I_{\frac{1}{3}}(\frac{2}{3}z^{\frac{3}{2}})\big),\quad z\in{\mathbb{C}},$
@@
 Then by (3.29), the zeros of ${\mathcal{C}}(z)$ in the sector $\big\{z\in{\mathbb{C}}|\ -\frac{2\pi}{3}\leq\text{Arg}(z)\leq\frac{\pi}{3}\big\}$ lie only on rays $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=-\frac{\pi}{3}\ \text{or}\ \frac{\pi}{3}\big\}$.
-Note that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big{\{}z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big{\}}$.
-By homogenization of (1.15), it is straight to obtain the linearized problem:
-(3.31) $\left\{\begin{aligned} &y\partial_{x}f-\partial_{y}^{2}f=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}f\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[f]=A+F,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
-In order to solve the problem (3.31), we first omit the integral condition.
-Lemma 3.2.
+Notice that ${\mathcal{C}}(z)$ is entire and ${\mathcal{C}}(z)={\mathcal{C}}(ze^{\frac{2\pi i}{3}})$ for any $z\in{\mathbb{C}}$, then we have that all zeros of ${\mathcal{C}}(z)$ in $\mathbb{C}$} belong to $\big\{z\in{\mathbb{C}}|\ \text{Arg}(z)=\frac{\pi}{3}+\frac{2k\pi}{3},\ k\in{\mathbb{Z}}\big\}$.
+We next solve the following homogeneous problem
+(3.30) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
+Proposition 3.2.
 Let $\bar{\alpha}\geq 0$ and $g=\partial_{x}|\partial_{x}|A\in H^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}\cap\dot{H}^{-1}$.
-Then, there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ of the boundary-value problem:
-(3.32) $\left\{\begin{aligned} &y\partial_{x}\omega_{b}-\partial_{y}^{2}\omega_{b}=0,\quad(x,y)\in\Omega_{M},\\ &\partial_{y}\omega_{b}\big{|}_{y=0}=g,\quad x\in{\mathbb{R}},\end{aligned}\right$.
-with
+Then there exists a solution $\omega_{b}\in X_{\bar{\alpha}}$ to the boundary-value problem (3.30) with the estimate
 (3.31) ${\big\|\omega_{b}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})},$
@@
 Proof.
-To solve the boundary-value problem (3.32), we first take the Fourier transform in variable $x$. Then we obtain
+To solve the boundary-value problem (3.30), we take the Fourier transform in variable $x$. Then we have
 (3.32) $\left\{\begin{aligned} i\xi y\hat{\omega}_{b}-\partial_{y}^{2}\hat{\omega}_{b}=&\ 0,\quad(\xi,y)\in{\mathbb{R}}\times(0,M)\\ \partial_{y}\hat{\omega}_{b}\big|_{y=0}=\hat{g}=&\ i\xi|\xi|\hat{A},\quad\xi\in{\mathbb{R}}.\end{aligned}\right$.
-The general solution to this BVP (3.34) is
+The general solution to the BVP (3.32) is as follows
 (3.33) $\hat{\omega}_{b}(\xi,y)=C(\xi)\bigg(\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg)+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}\hat{g}$.
-Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula as following:
+Taking $C(\xi)=0$, we obtain a solution $\omega_{b}$ with the explicit formula
 (3.34) $\hat{\omega}_{b}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}\hat{g}(\xi),\quad\xi<0.\end{aligned}\right$.
@@
 (3.35) ${\big\|(|\partial_{x}|^{\frac{1}{3}}y)^{\alpha^{\prime}}\omega_{b}\big\|}_{L^{2}(\Omega_{M})}$ $\leq C{\big\|(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\hat{\omega}_{b}\big\|}_{L^{2}(\Omega_{M})}$
-$\qquad\leq C{\big{\|}{\big{\|}Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big{\|}}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big{\|}}_{L_{\xi}^{2}}$
-$\qquad\leq C{\big{\|}|\xi|^{\frac{3}{2}}\hat{A}(\xi)\big{\|}}_{L_{\xi}^{2}}=C{\big{\|}|\partial_{x}|^{\frac{3}{2}}A\big{\|}}_{L^{2}({\mathbb{R}})}$
-$\qquad\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}A\big{\|}}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
-Since $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
-(3.38) ${\big{\|}f\big{\|}}_{Y_{\bar{\alpha}}}=$ $\ {\big{\|}y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}$
-$\ \ +{\big{\|}|\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}f\big{\|}}_{L_{x,y}^{2}}+{\big{\|}|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big{\|}}_{L_{x,y}^{2}}$
+$\leq C{\big\|{\big\|Ai((i\xi)^{\frac{1}{3}}y)(|\xi|^{\frac{1}{3}}y)^{\alpha^{\prime}}\big\|}_{L_{y}^{2}}\cdot|\xi|^{\frac{5}{3}}|\hat{A}(\xi)|\big\|}_{L_{\xi}^{2}}$
+$\leq C{\big\||\xi|^{\frac{3}{2}}\hat{A}(\xi)\big\|}_{L_{\xi}^{2}}=C{\big\||\partial_{x}|^{\frac{3}{2}}A\big\|}_{L^{2}({\mathbb{R}})}$
+$\leq C{\big\||\partial_{x}|^{\frac{5}{6}}A\big\|}_{H^{\frac{2}{3}}({\mathbb{R}})}.$
+As $g\in H^{\frac{1}{6}+\frac{\alpha}{2}}({\mathbb{R}})$, we obtain
+(3.36) ${\big\|\omega_{b}\big\|}_{Y_{\bar{\alpha}}}=$ $\ {\big\|y|\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{x}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{2}{3}+\frac{\bar{\alpha}}{2}}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{1+\bar{\alpha}}{2}}y^{\frac{1}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}$
+$\ \ +{\big\||\partial_{x}|^{\frac{1}{3}+\frac{\bar{\alpha}}{2}}\partial_{y}\omega_{b}\big\|}_{L_{x,y}^{2}}+{\big\||\partial_{x}|^{\frac{\bar{\alpha}}{2}}\partial_{y}^{2}\omega_{b}\big\|}_{L_{x,y}^{2}}$
 $\leq$ $\ C{\big\||\partial_{x}|^{\frac{1}{6}+\frac{\bar{\alpha}}{2}}g\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{13+3\bar{\alpha}}{6}}A\big\|}_{L^{2}({\mathbb{R}})}.$
-It follows $\omega_{b}\in X_{\bar{\alpha}}$ from (3.37) and (3.38). In particular, we have
-(3.39) ${\big{\|}\omega_{b}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
-Now, we are in the position to prove Theorem 1.1. Denote $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to
+It follows from (3.35) and (3.36) that $\omega_{b}\in X_{\bar{\alpha}}$ with
+${\big\|\omega_{b}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}.$
+Now we are in a position to address the system (3.1), i.e., complete the proof of Theorem 1.1. Let $(\omega_{0},u_{0},v_{0},A_{0})$ be the solution to the linear system
 (3.37) $\left\{\begin{aligned} &y\partial_{x}\omega_{0}-\partial_{y}^{2}\omega_{0}=0,\\ &u_{0}=I_{y}[\omega_{0}],\quad v_{0}=-\partial_{x}I_{y}[u_{0}],\\ &\partial_{y}\omega_{0}\big|_{y=0}=\partial_{x}|\partial_{x}|A_{0}(x),\quad I_{M}[\omega_{0}]=A_{0}(x)+F(x).\end{aligned}\right$.
@@
 Proof of Theorem 1.1.
-We hope that we take (3.36) as a solution to (3.40), that is,
+We take (3.34) as a solution to (3.37), that is,
 (3.38) $\hat{\omega}_{0}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}(i\xi|\xi|\hat{A}_{0})=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}}i\xi|\xi|^{\frac{2}{3}}\hat{A}_{0}(\xi),\quad\xi<0.\end{aligned}\right$.
@@
 Here
-(3.44) $p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
-Indeed, the integral is well-defined since $Ai(z)$ is analytic. We set
+$p_{M}(\xi)=$ $\ \int_{0}^{(i\xi)^{\frac{1}{3}}M}Ai(z)dz.$
+Indeed, the integral is well-defined, since $Ai(z)$ is analytic. Let
 (3.41) $m(\xi)=$ $1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}.$
-Then by (3.43), we obtain
+Then we get by (3.40) that
 (3.42) $\hat{A}_{0}=-\frac{1}{m(\xi)}\hat{F}$.
@@
 (3.43) $\hat{\omega}_{0}(\xi,y)=-\frac{Ai(e^{\frac{{\rm sgn\ }\xi}{6}\pi i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{6}\pi i}}\frac{i\xi|\xi|^{\frac{2}{3}}}{m(\xi)}\hat{F}(\xi).$
-To ensure that the formulas (3.46) and (3.47) are well-defined, we need $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. Then, we give an estimate of $m(\xi)$. We define
+To ensure that the formulas (3.42) and (3.43) are well-defined, we need to prove $m(\xi)\not=0$ for any $\xi\in{\mathbb{R}}$. We define
 $p_{0}:=\mathop{sup}\limits_{{\rm Arg\ }z\in\{\pm\frac{\pi}{6}\}}\bigg|\int_{0}^{z}Ai(\zeta)d\zeta\bigg|.$
-Note that $0\leq p_{0}<\infty$, then we obtain
+Noting that $0\leq p_{0}<\infty$, we obtain
 $a_{0}:=\bigg(\frac{4p_{0}}{3|Ai^{\prime}(0)|}\bigg)^{-\frac{3}{4}}>0.$
 It follows that
-(3.48) $m(\xi)\geq\frac{1}{4},\quad\forall\ |\xi|\leq a_{0}.$
+(3.44) $m(\xi)\geq\frac{1}{4}\quad\forall\ |\xi|\leq a_{0}.$
 Using the following identity
-(3.49) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3}.$
-As $M\to\infty$, the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$. By calculations, we have
+(3.45) $\int_{0}^{\infty}Ai(z)dz=\frac{1}{3},$
+the multiplier $m(\xi)$ converges to $\frac{1}{3}$ uniformly on $a_{0},\infty)$ as $M\to\infty$. A direct calculation yields
 (3.46) $\bigg|1-\frac{1}{3Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg|^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
-Then there exists a large enough constant $M_{0}>0$, when $M\geq M_{0}$, by (3.48) and (3.50), we also obtain
+Then there exists a large enough constant $M_{0}>0$ such that for $M\geq M_{0}$, we get by (3.44) and (3.46) that
 (3.47) $|m(\xi)|^{-1}=\bigg|1-\frac{p_{M}(\xi)}{Ai^{\prime}(0)e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}i\xi|\xi|^{\frac{1}{3}}\bigg|^{-1}\leq C(1+|\xi|)^{-\frac{4}{3}},\quad\forall\ \xi\in{\mathbb{R}}.$
-It concludes that $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, by Fourier multiplier theorem, there exists a constant $C>0$ such that
-(3.52) ${\big{\|}m(D)^{-1}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}f\big{\|}}_{L^{2}({\mathbb{R}})},\quad\forall\ f\in L^{2}({\mathbb{R}}).$
-Using Lemma 3.2 and the estimate (3.52), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as following:
-(3.53) ${\big{\|}\omega_{0}\big{\|}}_{X_{\bar{\alpha}}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big{\|}}_{L^{2}({\mathbb{R}})}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}F\big{\|}}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
-For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we obtain
-(3.54) ${\big{\|}u_{0}\big{\|}}_{L^{\infty}(\Omega_{M})}+{\big{\|}A_{0}\big{\|}}_{L^{\infty}({\mathbb{R}})}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big{\|}\omega_{0}\big{\|}}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big{\|}\omega_{0}\big{\|}}_{X_{0}}.$
-It completes the proof. ∎
+Thus, $\hat{A}_{0}$ and $\hat{\omega}_{0}$ are well-defined. Moreover, there exists a constant $C>0$ such that
+(3.48) ${\big\|m(D)^{-1}f\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\|f\big\|}_{L^{2}({\mathbb{R}})}.$
+Using Proposition 3.2 and (3.48), for any $\bar{\alpha}\geq 0$, we obtain the $X_{\bar{\alpha}}$-estimate of $\omega_{0}$ as follows
+${\big\|\omega_{0}\big\|}_{X_{\bar{\alpha}}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}A_{0}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\bar{\alpha}}{2}}F\big\|}_{L^{2}({\mathbb{R}})}\leq C{\big\||\partial_{x}|^{\frac{5}{6}}F\big\|}_{\dot{H}^{\frac{\bar{\alpha}}{2}}({\mathbb{R}})}.$
+For the bound of displacement $A_{0}$ and the velocity $u_{0}(x)$, we have
+${\big\|u_{0}\big\|}_{L^{\infty}(\Omega_{M})}+{\big\|A_{0}\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\omega_{0}\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big\|\omega_{0}\big\|}_{L_{y}^{1}L_{x}^{\infty}}\leq{\big\|\omega_{0}\big\|}_{X_{0}}.$
+This completes the proof. ∎
 Remark 3.2.
 Note that ${\rm Re\ }\bigg(\frac{p_{M}(\xi)}{e^{\frac{{\rm sgn\ }\xi}{3}\pi i}}\bigg)>0$ for $|\xi|^{\frac{1}{3}}M>0$ by numberical results (see figure 2, where $r=|\xi|^{\frac{1}{3}}M$).
-Then by continuity of $m(\xi)$, we have $\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}$ for any $\xi\in{\mathbb{R}}$ and $M>0$, where $C>0$ depends on $M^{-1}$.
-Figure 2. Analysis about $m(\xi)$.
-4. Proof of Main Theorem
-Let us recall the perturbation system (1.15) that
-(4.1) $\left\{\begin{aligned} &y\omega_{x}-\omega_{yy}=-u\omega_{x}-v\omega_{y},\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\omega_{y}\big{|}_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
-In order to find a solution of the equation (4.1), we use the iteration method. Set
+Then by the continuity of $m(\xi)$, we have
+$\frac{1}{|m(\xi)|}\leq C(1+|\xi|)^{-\frac{4}{3}}\text{ for any }\xi\in{\mathbb{R}}\text{ and }M>0,$
+where $C>0$ depends on $M^{-1}$.
+Figure 2. Analysis on $m(\xi)$.
+4. Existence and uniqueness of classical solution
+In this section, we establish the existence and uniqueness of classical solution to the nonlinear system (1.13), i.e.,
+(4.1) $\left\{\begin{aligned} &y\partial_{x}\omega-\partial_{y}^{2}\omega=-u\partial_{x}\omega-v\partial_{y}\omega,\\ &u=I_{y}[\omega],\quad v=-\partial_{x}I_{y}[u],\\ &\partial_{y}\omega\big|_{y=0}=\partial_{x}|\partial_{x}|A,\quad I_{M}[\omega]=A(x)+F(x).\end{aligned}\right$.
+To construct a solution to the system (4.1), we use the iteration method. Let
 $(\bar{\omega},\bar{A},\bar{u},\bar{v})=(\omega-\omega_{0},A-A_{0},u-u_{0},v-v_{0}).$
-Then, we reduce (4.1) to the following BVP:
+Then the system (4.1) is reduced to the following BVP:
 (4.2) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}-\partial_{y}^{2}\bar{\omega}=-(\bar{u}+u_{0})\partial_{x}(\bar{\omega}+\omega_{0})-(\bar{v}+v_{0})\partial_{y}(\bar{\omega}+\omega_{0}),\\ &\bar{u}=I_{y}[\bar{\omega}],\quad\bar{v}=-\partial_{x}I_{y}[\bar{u}],\\ &\partial_{y}\bar{\omega}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A},\quad I_{M}[\bar{\omega}]=\bar{A}.\end{aligned}\right$.
-Thus, we can consider the iterated system as following:
+We introduce the iterated system as follows
 (4.3) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{k+1}-\partial_{y}^{2}\bar{\omega}_{k+1}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}),\\ &\bar{u}_{k+1}=I_{y}[\bar{\omega}_{k+1}],\quad\bar{v}_{k+1}=-\partial_{x}I_{y}[\bar{u}_{k+1}],\\ &\bar{u}_{k}=I_{y}[\bar{\omega}_{k}],\quad\bar{v}_{k}=-\partial_{x}I_{y}[\bar{u}_{k}],\\ &\partial_{y}\bar{\omega}_{k+1}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\quad I_{M}[\bar{\omega}_{k+1}]=\bar{A}_{k+1}.\end{aligned}\right$.
-Recall that the space $X_{\alpha,\infty}$ is defined as
-(4.4) $X_{\alpha,\infty}=\bigg{\{}f:{\mathbb{R}}\to{\mathbb{R}}\big{|}\ \ {\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}f\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}}f\big{\|}}_{L^{2}({\mathbb{R}})}+{\big{\|}f\big{\|}}_{L^{\infty}({\mathbb{R}})}<\infty\bigg{\}}.$
-First of all, we give some results about the iterated system (4.3).
+First of all, we present the following uniform estimate for the iterated system (4.3).
 Proposition 4.1.
-Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$.
-Suppose $k$ is a non-negative integer and $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
-Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ of (4.3).
-Moreover, for $\varepsilon\in(0,\frac{1}{6}]$, we have
+Given $\alpha\in(\frac{2}{3},\frac{7}{3}]$, assume that $(\bar{\omega}_{k},\bar{A}_{k})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
+Then there exists a solution $(\bar{\omega}_{k+1},\bar{A}_{k+1})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ to (4.3), which holds that for $\varepsilon\in(0,\frac{1}{6}]$,
 (4.4) ${\big\|\bar{\omega}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}$
-$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\big{[}\big{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\big{)}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\big{]}.$
+$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
 Here $C=C(\alpha,\varepsilon)>0$ is a constant.
 Proof.
-Set $\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0})$, we decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
-(4.6) $\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big{|}_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big{|}_{y=M}=0,\end{aligned}\right$.
-$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big{|}_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1}.\end{aligned}\right$.
-with
+We decompose $\bar{\omega}_{k+1}=\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}$, where $\bar{\omega}_{e,k+1}$ and $\bar{\omega}_{b,k+1}$ satisfy
+$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{e,k+1}-\partial_{y}^{2}\bar{\omega}_{e,k+1}=\bar{h}_{k},\\ &\partial_{y}\bar{\omega}_{e,k+1}\big|_{y=0}=0,\quad\bar{\omega}_{e,k+1}\big|_{y=M}=0,\end{aligned}\right$.
+and
+$\left\{\begin{aligned} &y\partial_{x}\bar{\omega}_{b,k+1}-\partial_{y}^{2}\bar{\omega}_{b,k+1}=0,\\ &\partial_{y}\bar{\omega}_{b,k+1}\big|_{y=0}=\partial_{x}|\partial_{x}|\bar{A}_{k+1},\end{aligned}\right$.
+together with
 (4.5) $I_{M}[\bar{\omega}_{e,k+1}+\bar{\omega}_{b,k+1}]=\bar{A}_{k+1}.$
-By Lemma 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the following form
+Here
+$\bar{h}_{k}=-(\bar{u}_{k}+u_{0})\partial_{x}(\bar{\omega}_{k}+\omega_{0})-(\bar{v}_{k}+v_{0})\partial_{y}(\bar{\omega}_{k}+\omega_{0}).$
+By Proposition 3.1, we obtain the unique solution $\bar{\omega}_{e,k+1}$ with the form
 (4.6) $\bar{\omega}_{e,k+1}=$ $\ \pi s_{M}(i\xi)^{-\frac{1}{3}}\bigg\{\mu_{M}(\xi,y)\int_{0}^{y}\mu_{0}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz$
 $\qquad\qquad\qquad+\mu_{0}(\xi,y)\int_{y}^{M}\mu_{M}(\xi,z)\hat{\bar{h}}_{k}(\xi,z)dz\bigg\}.$
-By Lemma 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ as following:
-(4.9) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
+By Proposition 3.2, we can find a solution $\bar{\omega}_{b,k+1}$ with
+(4.7) $\hat{\omega}_{b,k+1}(\xi,y)=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),&xi<0.\end{aligned}\right$.
 In view of (4.5), we have
@@
 Here $m=m(\xi)$ is defined in (3.41).
-By Lemma 3.1, Theorem 1.1, and Proposition 2.3, we obtain
+By Proposition 3.1, Theorem 1.1 and Proposition 2.1, we obtain
 (4.9) ${\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}\leq$ $\ C\big((1+M^{2}){\big\|\bar{h}_{k}\big\|}_{L^{2}(\Omega_{M})}+{\big\||\partial_{x}|^{\frac{\alpha}{2}}\bar{h}_{k}\big\|}_{L^{2}(\Omega_{M})}\big)$
-$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{0}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}\partial_{x}A_{0}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
-$\qquad+{\big{\|}\partial_{x}A_{k}\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big{\|}\partial_{x}F\big{\|}}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2})$
-$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}).$
-By Lemma 3.2 and Proposition 2.4, we have
-(4.12) ${\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}={\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
-[…]
-$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}},$
-${\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}\leq C{\big{\|}|\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big{\|}|\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big{\|}}_{H^{\frac{4}{3}}({\mathbb{R}})}$
-$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-Combining (4.11) with (4.12), one gets
-(4.13) ${\big{\|}\bar{\omega}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{\alpha,\infty}}$
-$\qquad\leq{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{\omega}_{b,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
-$\qquad\leq C{\big{\|}\bar{\omega}_{e,k+1}\big{\|}}_{X_{0}\cap X_{\alpha}}$
-$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}\bigg{[}({\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}^{2}+{\big{\|}A_{k}\big{\|}}_{X_{\alpha,\infty}}^{2})+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\bigg{]}.$
-Then it completes the proof.
-To get the uniform estimates of the solutions, we use the following results which can be got the induction method.
+$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{0}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|\partial_{x}A_{0}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}$
+$\qquad+{\big\|\partial_{x}A_{k}\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}+{\big\|\partial_{x}F\big\|}_{L^{\infty}\cap\dot{H}^{\frac{\alpha}{2}}}^{2}\big)$
+$\leq$ $\ CM^{\frac{5+3\varepsilon}{2}}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\big).$
+By Proposition 3.2, Lemma 2.3 and (4.8), we have
+(4.10) $\begin{split}{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}&={\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\||\partial_{x}|^{\frac{5}{6}}I_{M}[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}I[\bar{\omega}_{e,k+1}]\big\|}_{L^{2}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\\ {\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}&\leq C{\big\||\partial_{x}|^{\frac{5}{6}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}+C{\big\||\partial_{x}|^{\frac{5}{6}+\frac{\alpha}{2}}\bar{A}_{k+1}\big\|}_{H^{\frac{4}{3}}({\mathbb{R}})}\\ &\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}},\end{split}$
+from which and (4.9), we arrive at
+${\big\|\bar{\omega}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{\alpha,\infty}}$
+$\leq{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{\omega}_{b,k+1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
+$\leq C{\big\|\bar{\omega}_{e,k+1}\big\|}_{X_{0}\cap X_{\alpha}}$
+$\leq CM^{\frac{5+3\varepsilon}{2}}\Big[\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}^{2}+{\big\|A_{k}\big\|}_{X_{\alpha,\infty}}^{2}\big)+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\Big].$
+The proof is completed. ∎
+To obtain the uniform estimates of the solution sequence, we use the following basic lemma, which can be proved by the induction.
 Lemma 4.1.
 Let $\alpha_{i},\beta_{i}\ (i=1,2)$ be positive constants. Assume that $\{a_{k}\}_{k\geq 0},\ \{b_{k}\}_{k\geq 1}$ are two non-negative sequences satisfying
-(4.14) $(1)\ a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2},\quad\mbox{for any }k\geq 0;\quad(2)\ b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2},\quad\mbox{for any }k\geq 1.$
-Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$, then we have
-(4.15) $0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
-Now, we are in the position to give the proof of Theorem 1.2.
+$a_{k+1}\leq\alpha_{1}a_{k}^{2}+\alpha_{2}\quad\mbox{for any }k\geq 0;\quad b_{k+1}\leq\beta_{1}b_{k}+\beta_{2}b_{k}^{2}\quad\mbox{for any }k\geq 1.$
+Assume that $\alpha_{2}\leq\frac{1}{4\alpha_{1}}$ and $\beta_{1}<1$. Set $a_{0}=0$ and $b_{1}\leq\frac{\gamma-\beta_{1}}{\beta_{2}}$ for some $\gamma\in(\beta_{1},1)$. Then we have
+$0\leq\sup_{k\geq 0}a_{k}\leq 2\alpha_{2},\qquad\sum_{k=1}^{\infty}b_{k}\leq\frac{b_{1}}{1-\gamma}.$
+With the above preparations, we complete the proof of Theorem 1.2.
 Proof of Theorem 1.2.
-The proof proceeds in three steps.
-Step 1. Bounded sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
+The proof consists of three steps.
+Step 1. Boundedness of the solution sequence.
 Taking $\bar{\omega}_{0}=\bar{A}_{0}=0$, by Proposition 4.1, we obtain a sequence $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$.
@@
 Clearly, we have $a_{0}=0$.
-By Proposition 4.1, one gets
-(4.16) $a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
-Using Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
-(4.17) $\sup_{k\geq 0}\bigg{(}{\big{\|}\bar{\omega}_{k}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\bar{A}_{k}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}^{2}.$
-Step 2. Convergence of this sequence.
+By Proposition 4.1, we have
+$a_{k+1}\leq CM^{\frac{5+3\varepsilon}{2}}a_{k}^{2}+CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
+By Lemma 4.1, for $4C^{2}M^{5+3\varepsilon}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}\leq 1$ (that is, ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$), we obtain
+$\sup_{k\geq 0}\big({\big\|\bar{\omega}_{k}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\bar{A}_{k}\big\|}_{X_{\alpha,\infty}}\big)=\sup_{k\geq 0}a_{k}\leq 2CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}^{2}.$
+Step 2. Convergence of the solution sequence.
 Set
@@
 […]
-We set
+We denote
 […]
@@
 […]
+By Lemma 4.1, if $CM^{\frac{5+3\varepsilon}{2}}{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2}$ and
+[…]
+which means that ${\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$, then there holds
 […]
 […]
-By Lemma 4.1, if $CM^{\frac{5+3\varepsilon}{2}}{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2}$ and
+Step 3. Uniqueness of the solution $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
+By Theorem 1.1, we know $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$.
+Combing (4.6)-(4.7) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we conclude that $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
+$\hat{\bar{\omega}}(\xi,y)\big|_{y=M}$ $=\hat{\bar{\omega}}_{b}(\xi,y)\big|_{y=M}$
+$=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0.\end{aligned}\right$.
+This shows that $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$.
+Moreover, it is easy to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}\setminus\{0\}$ due to $F\not\equiv 0$, which implies that the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial.
+Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions to the perturbation system (1.13). Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, which solve
 […]
-which means ${\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\leq\frac{1}{2CM^{\frac{5+3\varepsilon}{2}}}$ for $M\gg 1$. Then there holds
+As in the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
+$\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big|_{y=0}=0,\quad\tilde{\omega}_{e}\big|_{y=M}=0,\end{aligned}\right$.
+and
 […]
-[…]
-Step 3. Uniqueness of $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})$ in ${\mathcal{X}}_{\alpha,M}$.
-Recall that
-${\mathcal{X}}_{\alpha,M}$= $\ \bigg{\{}(\omega,A)\big{|}\ {\big{\|}(\omega,A)\big{\|}}_{{\mathcal{X}}_{\alpha,M}}={\big{\|}\omega\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}A\big{\|}}_{X_{\alpha,\infty}}<\infty,$
-$\qquad\qquad\quad\omega(\cdot,y)\big{|}_{y=M}=\frac{Ai(M\partial_{x}^{\frac{1}{3}})}{Ai^{\prime}(0)}\partial_{x}^{\frac{2}{3}}|\partial_{x}|A\bigg{\}}.$
-By Theorem 1.1, we obtain $(\omega_{0},A_{0})\in{\mathcal{X}}_{\alpha,M}$.
-Combing (4.8)-(4.9) with the convergence of $\{(\bar{\omega}_{k},\bar{A}_{k})\}_{k\geq 0}$ in $X_{0}\cap X_{\alpha}\times X_{\alpha,\infty}$, we have $(\bar{\omega},\bar{A})\in(X_{0}\cap X_{\alpha})\times X_{\alpha,\infty}$ and
-(4.22) $\hat{\bar{\omega}}(\xi,y)\big{|}_{y=M}=\hat{\bar{\omega}}_{b}(\xi,y)\big{|}_{y=M}$
-$\qquad=\left\{\begin{aligned} &\frac{Ai(e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{\frac{\pi}{6}i}\xi^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi>0,\\ &\frac{Ai(e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)e^{-\frac{\pi}{6}i}(-\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\bar{A}}_{k+1}(\xi),\quad\xi<0,\end{aligned}\right$.
-which implies $(\bar{\omega},\bar{A})\in{\mathcal{X}}_{\alpha,M}$.
-Moreover, it is evident to verify that $(\omega_{0}+\bar{\omega},A_{0}+\bar{A})\in{\mathcal{X}}_{\alpha,M}-\{0\}$ due to $F\not\equiv 0$, which implies the space ${\mathcal{X}}_{\alpha,M}$ is non-trivial.
-Let $(\omega_{i},A_{i})\ (i=1,2)\in{\mathcal{X}}_{\alpha,M}$ be two solutions of the perturbation system (1.15).
-Set $\tilde{\omega}=\omega_{1}-\omega_{2},\ \tilde{A}=A_{1}-A_{2}$, then there holds
-[…]
-Similar to the proof of Proposition 4.1, we decompose $\tilde{\omega}=\tilde{\omega}_{e}+\tilde{\omega}_{b}$ with
-(4.24) $\left\{\begin{aligned} &y\partial_{x}\tilde{\omega}_{e}-\partial_{y}^{2}\tilde{\omega}_{e}=\tilde{h},\\ &\partial_{y}\tilde{\omega}_{e}\big{|}_{y=0}=0,\quad\tilde{\omega}_{e}\big{|}_{y=M}=0,\end{aligned}\right$.
-[…]
-By Lemma 3.1, we have
+By Proposition 3.1, we have
 (4.13) ${\big\|\tilde{\omega}_{e}\big\|}_{X_{0}\cap X_{\alpha}}\leq$ $\ CM^{2}{\big\|\tilde{h}\big\|}_{L^{2}(\Omega_{M})}+C{\big\|\tilde{h}\big\|}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}.$
-By Lemma 3.2, we obtain
-(4.26) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\bigg{(}\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\bigg{)}+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
-In view of $\hat{\tilde{\omega}}_{e}\big{|}_{y=M}=0$, one gets
-(4.27) $\hat{\tilde{\omega}}_{b}\big{|}_{y=M}=\hat{\tilde{\omega}}\big{|}_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$.
-Combining (4.26) with (4.27), we obtain
+By Proposition 3.2, we get
+(4.14) $\hat{\tilde{\omega}}_{b}(\xi,y)=C(\xi)\big(\sqrt{3}Ai((i\xi)^{\frac{1}{3}}y)+Bi((i\xi)^{\frac{1}{3}}y)\big)+\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
+In view of $\hat{\tilde{\omega}}_{e}\big|_{y=M}=0$, we have
+$\hat{\tilde{\omega}}_{b}\big|_{y=M}=\hat{\tilde{\omega}}\big|_{y=M}=\frac{Ai((i\xi)^{\frac{1}{3}}M)}{Ai^{\prime}(0)}(i\xi)^{\frac{2}{3}}|\xi|\hat{\tilde{A}}$,
+which along with (4.14) gives
 $C(\xi)=0.$
 Thus, we have
-(4.28) $\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
-This formula is identical to (3.36), then using the Lemma 3.2, we obtain
+$\hat{\tilde{\omega}}_{b}(\xi,y)=\frac{Ai((i\xi)^{\frac{1}{3}}y)}{Ai^{\prime}(0)(i\xi)^{\frac{1}{3}}}i\xi|\xi|\hat{\tilde{A}}$.
+This formula is identical to (3.34). We then get by Proposition 3.2 that
 (4.15) ${\big\|\tilde{\omega}_{b}\big\|}_{X_{0}\cap X_{\alpha}}\leq C{\big\||\partial_{x}|^{\frac{5}{6}}\tilde{A}\big\|}_{H^{\frac{4}{3}+\frac{\alpha}{2}}}\leq C{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}.$
-Since $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
-(4.30) $\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
-It follows from Proposition 2.4 that
+Thanks to $I_{M}[\tilde{\omega}_{e}+\tilde{\omega}_{b}]=\tilde{A}$, we have
+$\tilde{A}=m(-i\partial_{x})^{-1}I_{M}[\tilde{\omega}_{e}].$
+It follows from Lemma 2.3 that
 (4.16) ${\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\leq C{\big\|\tilde{\omega}_{e}\big\|}_{X_{0}\cap X_{\alpha}}.$
 By Remark 2.1, we obtain
-(4.32) ${\big{\|}\tilde{h}\big{\|}}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
-$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}},$
-${\big{\|}\tilde{h}\big{\|}}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\bigg{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\bigg{)}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}$
-$\qquad\qquad\quad+CM\bigg{(}{\big{\|}\tilde{\omega}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}\bigg{)}{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}.$
-On the other hand, we know
+(4.17) ${\big\|\tilde{h}\big\|}_{L^{2}}\leq\ CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
+$\qquad\qquad\quad+CM^{\frac{1+3\varepsilon}{2}}\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}},$
+${\big\|\tilde{h}\big\|}_{\dot{H}_{x}^{\frac{\alpha}{2}}L_{y}^{2}}\leq CM\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big){\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}$
+$\qquad\qquad\quad+CM\big({\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}\big){\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}.$
+On the other hand, we know that
 (4.18) ${\big\|\tilde{A}\big\|}_{L^{\infty}({\mathbb{R}})}\leq{\big\|\tilde{\omega}\big\|}_{L_{x}^{\infty}L_{y}^{1}}\leq{\big\|\tilde{\omega}\big\|}_{X_{0}\cap X_{\alpha}}.$
-Combining (4.25), (4.29), (4.31), (4.32) with (4.33), we obtain
-(4.34) ${\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
-$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
-$\qquad\qquad\cdot\big{(}{\big{\|}\omega_{1}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\omega_{2}\big{\|}}_{X_{0}\cap X_{\alpha}}+{\big{\|}\tilde{A}\big{\|}}_{X_{\alpha,\infty}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}$
-$\qquad\leq CM^{\frac{5+3\varepsilon}{2}}{\big{\|}(\tilde{\omega},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}$
-$\qquad\qquad\cdot\big{(}{\big{\|}(\omega_{1},\tilde{A})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}(\omega_{2},A_{2})\big{\|}}_{{\mathcal{X}}_{\alpha,M}}+{\big{\|}F\big{\|}}_{H^{2}({\mathbb{R}})}\big{)}.$
+Summing up (4.13), (4.15), (4.16), (4.17) and (4.18), we obtain
+${\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}$
+$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|\omega_{1}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\omega_{2}\big\|}_{X_{0}\cap X_{\alpha}}+{\big\|\tilde{A}\big\|}_{X_{\alpha,\infty}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big)$
+$\leq CM^{\frac{5+3\varepsilon}{2}}{\big\|(\tilde{\omega},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}\big({\big\|(\omega_{1},\tilde{A})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|(\omega_{2},A_{2})\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\big).$
 This implies the uniqueness of solution $(\omega,A)$ in ${\mathcal{X}}_{\alpha,M}$ to (1.13) when ${\big\|(\omega,A)\big\|}_{{\mathcal{X}}_{\alpha,M}}+{\big\|F\big\|}_{H^{2}({\mathbb{R}})}\leq\frac{\delta_{0}}{M^{\frac{5+3\varepsilon}{2}}}$ with $\delta_{0}=\delta_{0}(\alpha,\varepsilon)\ll 1$.
-Appendix
+∎
+Appendix A The physical derivation of triple-deck equation
 This appendix provides a brief overview of the triple-deck formalism, based on the physical model of an incompressible boundary layer with a roughness element; for further details, see [26].
@@
 The system is described in the Cartesian coordinate $(X,Y)$, with its origin at the leading edge.
-The roughness height is assumed to be $O(\epsilon^{5})$ or smaller, and its streamwise length scale is $O(\epsilon^{3})$.
-Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the $O(\epsilon^{3})$ vicinity of the roughness (near $X=1$).
-The three distinct layers are summarized as follows:
-I. Viscous sublayer: Vertical scale $O(\epsilon^{5})$, streamwise scale $O(\epsilon^{3})$.
-II. Main layer: Vertical scale $O(\epsilon^{4})$, streamwise scale $O(\epsilon^{3})$.
-III. Upper layer: Vertical scale $O(\epsilon^{3})$, streamwise scale $O(\epsilon^{3})$.
+The roughness height is assumed to be ${\mathcal{O}}(\epsilon^{5})$ or smaller, and its streamwise length scale is ${\mathcal{O}}(\epsilon^{3})$.
+Consequently, the roughness-induced mean-flow distortion exhibits three distinct layers in the transverse direction, spanning in the ${\mathcal{O}}(\epsilon^{3})$ vicinity of the roughness (near $X=1$).
+The three distinct layers are summarized as follows.
+I. Viscous sublayer: Vertical scale ${\mathcal{O}}(\epsilon^{5})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
+II. Main layer: Vertical scale ${\mathcal{O}}(\epsilon^{4})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
+III. Upper layer: Vertical scale ${\mathcal{O}}(\epsilon^{3})$, streamwise scale ${\mathcal{O}}(\epsilon^{3})$.
 These scales arise from matched asymptotic expansions, connecting the triple-deck theory to classical boundary layer theory in the outer flow region.
@@
 In the main layer (layer II), we introduce a pair of local coordinates,
-$x_{*}=\epsilon^{-3}(X-1)=O(1),\quad\tilde{y}=\epsilon^{-4}Y=O(1).$
-The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+O(\epsilon^{4})$ and a mean-flow distortion,
+$x_{*}=\epsilon^{-3}(X-1)={\mathcal{O}}(1),\quad\tilde{y}=\epsilon^{-4}Y={\mathcal{O}}(1).$
+The flow field in this layer can be expressed as a sum of the Blasius base flow ${\mathbf{u}}=(\tilde{U}_{0}(y),0)+{\mathcal{O}}(\epsilon^{4})$ and a mean-flow distortion,
 (A.2) $u_{NS}(X,Y;\epsilon)=$ $\ \tilde{U}_{0}(\tilde{y})+\epsilon\tilde{u}(x_{*},\tilde{y})+o(\epsilon),$
@@
 (A.3) $\left\{\begin{aligned} \tilde{U}_{0}\partial_{x_{*}}\tilde{u}+\tilde{U}_{0}^{\prime}\tilde{v}=&\ 0,\\ \partial_{\tilde{y}}\tilde{p}=&\ 0,\\ \partial_{x_{*}}\tilde{u}+\partial_{\tilde{y}}\tilde{v}=&\ 0.\end{aligned}\right$.
-The solutions of this system are
+The solutions to this system are
 (A.4) $\tilde{u}=\lambda^{-1}A_{*}(x_{*})\tilde{U}_{0}^{\prime}(\tilde{y}),\quad\tilde{v}=-\lambda^{-1}A_{*}^{\prime}(x_{*})\tilde{U}_{0}(\tilde{y}),\quad\tilde{p}=\tilde{p}(x_{*}).$
@@
 In this upper layer, we introduce the local coordinate
-$\bar{y}=\epsilon^{-3}Y=O(1).$
+$\bar{y}=\epsilon^{-3}Y={\mathcal{O}}(1).$
 Matching with the upper limit of the main-layer solutions, we can express the flow field as
@@
 For convenience, we introduce
-$y_{*}=\epsilon^{-5}Y=O(1).$
-Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+O(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$.
+$y_{*}=\epsilon^{-5}Y={\mathcal{O}}(1).$
+Analysis of the Blasius equation leads to $\tilde{U}_{0}\to\lambda\tilde{y}+{\mathcal{O}}(\tilde{y}^{4})$ as $\tilde{y}\to 0^{+}$.
 Thus, the main-layer solution behaves like $u_{NS}=\bar{U}_{0}+\epsilon\tilde{u}+o(\epsilon)\to\epsilon(\lambda y_{*}+A_{*})+o(\epsilon)$, which determines the magnitude of the sublayer-layer solution.
@@
 The upper boundary condition of the sublayer equations should be obtained by matching with the main layer.
-Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}=O(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$.
+Now, we consider the overlapping region connecting the sublayer and main layer, where $y_{*}={\mathcal{O}}(\epsilon^{-\alpha_{0}})$ with $\alpha_{0}\in(0,1)$.
 Then for any $\beta_{0}>0$, we have
@@
 (A.16) $\big(u_{*}(x_{*},y_{*})-\lambda y_{*}\big)\big|_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad\tilde{p}(x_{*})=P_{*}(x_{*}).$
-From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as following:
+From (A.12)-(A.14) and (A.16), we can derive the classical complete system in viscous sublayer as follows
 (A.17) $\left\{\begin{aligned} &u_{*}\partial_{x_{*}}u_{*}+v_{*}\partial_{y_{*}}u_{*}+\frac{dP_{*}}{dx_{*}}-\partial_{y_{*}}^{2}u_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &\partial_{x_{*}}u_{*}+\partial_{y_{*}}v_{*}=0,\quad(x_{*},y_{*})\in\Omega_{\infty,F},\\ &[u_{*},v_{*}]\big|_{y_{*}=F(x_{*})}=0,\quad(u_{*}-\lambda y_{*})\big|_{y_{*}\uparrow\infty}=A_{*}(x_{*}),\quad x\in{\mathbb{R}},\\ &(u_{*}-\lambda y_{*})\big|_{x_{*}\downarrow-\infty}=0,\quad y_{*}\in{\mathbb{R}}_{+},\\ &P_{*}(x_{*})=-\frac{1}{\pi\lambda}\text{P.V.}\int_{{\mathbb{R}}}\frac{A_{*}^{\prime}(\zeta)}{\zeta-x_{*}}d\zeta,\quad x\in{\mathbb{R}}.\end{aligned}\right$.
@@
 Indeed, the above nonlinear system is usually solved using numerical approach.
-In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in O(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$.
+In light of (A.15), it is permissible to modify the upper boundary of the domain $\Omega_{\infty,F}$ to be a curve $\gamma=\{(x_{*},y_{*})|\ y_{*}=f(x_{*})\in{\mathcal{O}}(\epsilon^{-\alpha_{0}}),\ x_{*}\in{\mathbb{R}}\}$.
 Consequently, the solution $\phi_{*}=(u_{*},v_{*},P_{*})$ should be denoted by $\phi_{*,\epsilon}$ to include the impact of the upper boundary.
@@
 Typically, we take $\lambda=1$ in this paper.
-Acknowledgments.
+Acknowledgments
 M.
-Dong is partially supported by the Chinese Academy of Sciences (CAS) Project for Young Scientists in Basic Research under Grant No.
+Dong is partially supported by the NSFC under Grant No.
+12588201 and the CAS Project for Young Scientists in Basic Research under Grant No.
 YSBR-087.
 C.
-Wang is partially supported by the National Natural Science Foundation of China (NSFC) under Grant No.
+Wang is partially supported by the NSFC under Grant No.
 12471189.
@@
 12288101.
+Data availability The manuscript has no associated data.
+Declarations
+Conflict of interest The authors state that there is no conflict of interest.